### v1

# Discrete quantum groups and their duals

Abstract
Discrete quantum groups were introduced as duals of compact quantum groups by Podleś and Woronowicz in 1990 [Po-Wo]. Shortly after, they were defined and studied intrinsically in [Ef-Ru] and [VD-discrete]. In 1998, with the introduction of the multiplier Hopf algebras with integrals (also called algebraic quantum groups), the duality between discrete and compact quantum groups became part of the more general duality in the self-dual category of these algebraic quantum groups [VD-alg]. Again a few years later the duality was extended to all locally compact quantum groups (see e.g. [Ku-Va2]).
In these notes, we give a new and a somewhat updated approach of the theory of discrete quantum groups. In particular, we view them as special cases of algebraic quantum groups. The duality between the compact quantum groups and the discrete quantum groups is seen in this larger context. This has a number of advantages as we will explain.
On the one hand, we provide quite a bit of information about how all of this fits into the more general theory of algebraic quantum groups and its duality. Occasionally, we even go one step further and look at the most general case of locally compact quantum groups. Also sometimes, we compare with known results in pure Hopf algebra theory. On the other hand however, we have tried to make these notes highly self-contained. The aim of these notes in the first place is not to give new results but rather to review known results in a more modern perspective, taking into account recent developments. We believe this may be helpful for people who want to work with compact and discrete quantum groups now.
Date: 13 December 2025
    Department of Mathematics, KU Leuven, Celestijnenlaan 200B, B-3001 Heverlee, Belgium. E-mail: Alfons.VanDaele@kuleuven.be
0 . Introduction
The operator algebra approach to quantum groups finds it origin in the attempts to generalize Pontryagin’s duality theorem for abelian locally compact groups ([Po]) to the non-abelian ones. The first results in this direction all describe a dual of a (not necessarily abelian) locally compact group and show how to recover the original group from this dual object. Unfortunately, the nice symmetry, present in the abelian case, had to be given up. With the concept of a Kac algebra, developed independently by Kac and Vainerman ([Va-Ka]) and Enock and Schwarz ([En-Sw1, En-Sw2]), the duality was again realized within a sefl-dual category of objects.
However, the work of Drinfel’d ([Dr]) and Jimbo ([Ji]) made people realize that one of the conditions, imposed on the antipode in the theory of Kac algebras, was too restrictive. At the same time, Woronowicz ([Wo1]) constructed the quantum $SU_{q}(2)$ as a compact quantum group within the operator algebra framework. Also this example did not satisfy the Kac algebra axioms. These two events were the starting point for a new search for a self-dual category of objects, strictly larger than the category of Kac algebras so as to include also the compact quantum groups like Woronowicz’ $SU_{q}(2)$.
The result is now known as the theory of locally compact quantum groups. The development of the final object is due to Masuda, Nakagami and Woronowicz on the one hand ([Ma-Na, Ma-Na-Wo]) and Kustermans and Vaes on the other hand ([Ku-Va1, Ku-Va2, Ku-Va3], see also [VD-kerala, VD-sigma].
An intermediate step towards this result came with the theory of multiplier Hopf -algebras with positive integrals (sometimes called -algebraic quantum groups) ([VD-mha, VD-alg]). It is a category of objects, containing both the discrete and the compact quantum groups and it is also self-dual. It is not general enough to contain all locally compact quantum groups. On the other hand however, it is relatively simple and purely algebraic and the theory is rich enough to incorporate most algebraic features of the general theory, avoiding the more difficult technicalities.
The theory of locally compact quantum groups is now well-developed. It is well-understood and there are interesting and plenty of examples, and there are the common general construction methods. And although there is now also a simplified approach to the theory ([VD-sigma]), it remains technically quite involved and relies heavily on Tomita-Takesaki theory techniques. This is probably one of the reasons why it seems that presently, the theory is not very well absorbed by the mathematical community. Perhaps also the fact that there are so many differences in conventions used by various authors (see the appendix).
Lately there came a renewed interest in various aspects of the theory of discrete and compact quantum groups, but generally speaking, the main references that are used date from the period when the locally compact quantum groups, or even the algebraic quantum groups did not yet existed. On the other hand, there are good reasons to believe that a more modern point of view, where compact and discrete quantum groups are seen as special cases of locally compact quantum groups, in particular as cases of multiplier Hopf -algebras with positive integrals and where the duality between these two is viewed as a special case of the more general duality, has a lot of advantages. In particular, often the more general approach, certainly within the framework of the algebraic quantum groups, is simpler. But also, one might expect that the recent understanding of locally compact quantum groups can possibly contribute to a better way to treat discrete quantum groups and their duals.
These considerations take us to the aim and the content of these notes.
In the first place, we give a somewhat new and updated approach of the theory of discrete quantum groups. As mentioned already, discrete quantum groups were first introduced by Podleś and Woronowicz as duals of compact quantum groups ([Po-Wo]). Later, discrete quantum groups were intrinsically defined and studied by Effros and Ruan ([Ef-Ru]) and in [VD-discrete]. Therefore, we now look at the work of Podleś and Woronowicz as a way to show that the dual of a compact quantum group (or rather a compact quantum matrix group) is a discrete quantum group. In this sense, their result is a special case of the more general duality between locally compact quantum groups, in particular, between -algebraic quantum groups. Here, undoubtedly, the general case is simpler than the special case as it is treated in the paper by Podleś and Woronowicz. The extra property that the underlying algebra of the dual is a direct sum of matrix algebras follows in the first place from results within the theory of compact quantum groups, but it can also be obtained from the existence of the cointegral in the dual.
It should be mentioned that viewing discrete quantum groups as duals of compact quantum groups is mainly motivated by historical arguments. The notion of a discrete quantum group, obtained as a quantization of a discrete group, following the philosophy of quantization, is a very natural object. And in fact, the development of the theory is easier than that of compact quantum groups. We will comment more on this in the appendix.
We have made the updating of this approach to discrete quantum groups to a great extend in a self-contained way. However, as we want to emphasize also that the theory is part of a bigger context, we will regularly indicate how the results obtained are special cases of the general results. We will look here in the first place to the theory of algebraic quantum groups. Occasionally, we wil go further and consider the more general case of locally compact quantum groups. Also, when appropriate, we will compare with results in pure Hopf algebra theory.
The reader who has some background in these other and more general theories will benefit from these comments as he will be able to situate the results in a bigger context. The reader with no background of this kind may become anxious to learn more in this direction. He may even profit from these notes as they treat discrete and compact quantum groups more in the spirit of the general theory. However, it is possible to read and profit from this approach if none of this is the case. And we hope, all readers will enjoy the beauty (and simplicity) of this theory. After all, the main intention for distributing these notes is to make this material easily accessible.
Content of the paper
In Section 1, we start with a more general concept than the one of a discrete quantum group. We take a multiplier Hopf algebra $(A,\Delta)$ and we call it of discrete type if there is a left and a right cointegral (see Definitions 1.1 and 1.3). We are particularly interested in the case where a left cointegral $h$ is also a right cointegral. In that case, the element $\Delta(h)$, obtained by applying the coproduct on the left cointegral, plays a very special role. It is a separability idempotent in the multiplier algebra $M(A\otimes A)$. It is used to construct a left and a right integral on $A$ and many properties of these integrals are obtained from the properties of the cointegral.
In Section 2 we apply the results obtained in the first section to the special case of a discrete quantum group. This is defined as a multiplier Hopf -algebra $(A,\Delta)$ of discrete type with the extra constraint that $A$ is a direct sum of full matrix algebras $A_{\alpha}$ over $\mathbb{C}$, see Definition 2.2. In this case, the left cointegral can be constructed from the axioms of a multiplier Hopf -algebra whose underlying -algebra is such a direct sum. One of the main results here is that the square $S^{2}$ of the antipode $S$ leaves the components invariant. It is a homomorphism and has to be implemented by and invertible element on each component $A_{\alpha}$. As it turns out, it is in fact implemented by a positive invertible element $q_{\alpha}$ in $A_{\alpha}$. The integrals on $A$ are expressed in terms of the standard traces on the components, using these elements $q_{\alpha}$. Hence, we see that the integrals and the corresponding data, are all expressible using the data of the separability idempotent $\Delta(h)$. This treatment of discrete quantum groups is somewhat different from the original way as it is done in [VD-discrete].
In Section 3, we consider duality. In the first place, we look at the dual of a discrete quantum group which is a compact quantum group. The duality is seen within the framework of multiplier Hopf algebras with integrals. Then, we show the one-to-one correspondence between representations of the underlying algebra of the discrete quantum group and the corepresentations of the dual compact quantum group. We obtain various properties of compact quantum groups in this way. Observe once more that this approach is different from what is the common practice where first compact quantum groups are studied and then information about the dual is obtained. Our approach in these notes is justified as, after all, we basically study discrete quantum groups.
In Section 4, we simply discuss the classical cases of a discrete group and of the dual of a compact group. This is just to illustrate the preceding treatment of discrete quantum groups. In a separate paper [VD-suq2] we give a detailed treatment of the discrete quantum group $su_{q}(2)$ as another example of the theory of discrete quantum groups as it is studied in this paper. I believe that this is new although the underlying ideas are not. Indeed, in earlier work the discrete quantum group $su_{q}(2)$ has been studied as the dual of the compact quantum group $SU_{q}(2)$ as introduced by Woronowicz in [Wo1], see again [Po-Wo]. In [VD-suq2] we reverse the order and start with the discrete quantum group. The dual is then the compact quantum group $SU_{q}(2)$.
In the last section, Section 5 we draw some conclusions and discuss possible further research.
Finally, in an Appendix, we treat some more aspects of the differences in approaches to discrete and compact quantum groups, as well as the different conventions that are used by different authors.
Notations and conventions
Except for Section 1, we only work with -algebras over the complex numbers. The identity element in an algebra is denoted by $1$ while the identity in a group will be denoted by $e$. We use $\iota$ to denote the identity map. The linear dual of a vector space $V$ will be denoted by $V^{\prime}$.
In most cases, the tensor product is purely algebraic. If this is not the case, we will say so explicitly.
We use the leg-numbering notation for elements in, as well as operators on a tensor product. If e.g. $x$ is an element in the tensor product $A\otimes A$ of an algebra $A$ (with unit) with itself, then $x_{12}$, $x_{13}$ and $x_{23}$ are the elements in $A\otimes A\otimes A$ where $x$ is considered as sitting in respectively the first two factors, the first and the third factor and the second and the third factor. At the remaining place, we put the identity $1$ of the algebra. So, we get e.g. $x_{12}=x\otimes 1$. Similarly for linear operators where at the remaining factor, the identity map $\iota$ is placed.
Because it is a very convenient tool, sometimes in these notes, we will also use the Sweedler notation when $(A,\Delta)$ is a multiplier Hopf algebra. We write
$\Delta(a)=\sum_{(a)}a_{(1)}\otimes a_{(2)}\qquad\text{and}\qquad(\Delta\otimes\iota)\Delta(a)=\sum_{(a)}a_{(1)}\otimes a_{(2)}\otimes a_{(3)}$ (0.1)
when $a$ is an element of $A$. This can be done, provided we multiply in all but one factor with elements from $A$, left or right. The use of the Sweedler notation for multiplier Hopf algebras has been introduced in [Dr-VD]. The technique has been refined in [VD-tools] and in an appendix of [VD-lnalg]. For these discrete quantum groups, it is relatively transparent as one can simply multiply with central projections, in other words, one can restrict to components.
Basic references
The standard references for the theory of Hopf algebras are [Ab] and [Sw]. There is also the more recent work of Radford [Ra1].
The main references for multiplier Hopf algebras and algebraic quantum groups are [VD-mha] and [VD-alg]. Other possible references here are [VD-Zh1] and [VD-part1, VD-part2, VD-part3], as well as [Ti]. See also the more recent paper [DC-VD] with some new results. For discrete quantum groups we refer to [Po-Wo]], [Ef-Ru] and [VD-discrete] but our treatment here will be closest to [VD-discrete]. For compact quantum groups we have the basic references [Wo2] and [Wo3], see also [Ma-VD] for an other treatment.
The main references for the theory of locally compact quantum groups that we use are [Ku-Va2] and [Ku-Va3]. See also the newer and simplified approach in [VD-sigma].
Acknowledgments
I am very grateful to my colleagues and friends, for the warm hospitality while staying at the University of Fukuoka (Japan) where this work was initiated. I am also indebted to the KU Leuven for the opportunity to continue my research after my official retirement.
1 . Multiplier Hopf algebras of discrete type
We start this paper on discrete quantum groups with a more general concept.
$1.$
  • Definition Let $(A,\Delta)$ be a multiplier Hopf algebra. A non-zero element $h$ in $A$ satisfying $ah=\varepsilon(a)h$ for all $a\in A$ is called a left cointegral. A non-zero element $k$ in $A$ satisfying $ka=\varepsilon(a)k$ for all $a\in A$ is called a right cointegral.
First we have the following easy observation.
1.2.
  • Proposition If $h$ is a left cointegral satisfying $\varepsilon(h)\neq 0$, then it is also a right cointegral. It can be normalized such that $h^{2}=h$ and then it is unique.
  • Proof: If $h$ satisfies $\varepsilon(h)\neq 0$, we can take a scalar multiple and assume that $\varepsilon(h)=1$. Then $h^{2}=h$. Moreover, if $k$ is a right cointegral we have $kh=\varepsilon(h)k=\varepsilon(k)h$. So $k=\varepsilon(k)h$. Because $k$ is also non-zero we must have $\varepsilon(k)\neq 0$, we can also scale $k$ so that $\varepsilon(k)=1$ and then $h=k$.
    If $h^{\prime}$ is any other left cointegral, we have $hh^{\prime}=\varepsilon(h)h^{\prime}=h^{\prime}$ and because $h$ is a right cointegral we have $hh^{\prime}=\varepsilon(h^{\prime})h$. It follows that $h^{\prime}=\varepsilon(h^{\prime})h$ so that $h^{\prime}$ is a scalar multiple of $h$. Because $h^{\prime}$ is also assumed to be non-zero, we must have $\varepsilon(h^{\prime})\neq 0$ and we can scale $h^{\prime}$ so that $\varepsilon(h^{\prime})=1$. Then $h^{\prime}=h$ and we get uniqueness.
It can happen that $\varepsilon(h)=0$ and that $h$ and $k$ are different. Still, also in that case, one can show that cointegrals are unique, see [VD-Zh0].
If $(A,\Delta)$ is a regular multiplier Hopf algebra and $h$ is a left cointegral, then $S(h)$ is a right cointegral because for a regular multiplier Hopf algebra, the antipode is a bijective anti-isomorphism of $A$. If $\varepsilon(h)\neq 0$ we will have $S(h)=h$. But also in this case, it can happen that $\varepsilon(h)=0$. In fact that is already possible for finite-dimensional Hopf algebras.
If $(A,\Delta)$ is a multiplier Hopf -algebra and if $h$ is a left cointegral, we have $h^{*}h=\varepsilon(h^{*})h=\varepsilon(h)h$. If $\varepsilon(h)=0$ we would have $h^{*}h=0$. In a sense, this means that we have an involution that is not very useful. On the other hand, if $\varepsilon(h)\neq 0$ we see that we can scale $h$ so as to become a self-adjoint idempotent. Therefore, it makes sense to require that we have -algebras with the property that $a^{*}a=0$ implies that $a=0$. This is the case when $A$ is an operator algebra. We will impose this condition for a discrete quantum group in Section 2.
Further in this section, we will consider these different cases.
1.3.
  • Definition A multiplier Hopf algebra is called of discrete type if there is a left and a right cointegral.
We will now prove several properties of cointegrals. Most of them can be found already in [VD-Zh0].
1.4.
  • Proposition For a left cointegral $h$ and a right cointegral $k$ we have
    $\Delta(a)(1\otimes h)=a\otimes h\qquad\quad\text{and}\qquad\quad(k\otimes 1)\Delta(a)=k\otimes a$
    for all $a\in A$.
This follows immediately from the definitions.
Also the following result is still easy to obtain.
1.5.
  • Proposition Let $(A,\Delta)$ be any multiplier Hopf algebra. Assume that $h$ is a left cointegral in $A$. Then $(1\otimes a)\Delta(h)=(S(a)\otimes 1)\Delta(h)$ for all $a\in A$. If $k$ is a right cointegral, we have $\Delta(k)(a\otimes 1)=\Delta(k)(1\otimes S(a))$ for all $a$.
  • Proof: First remark the following. These equations hold in the multiplier algebra $M(A\otimes A)$, except when $(A,\Delta)$ is regular. Then they hold in $A\otimes A$.
    Using the Sweedler notation we write $1\otimes a=\sum_{(a)}(S(a_{(1)})\otimes 1)\Delta(a_{(2)})$. Then, for all $a\in A$,
    $(1\otimes a)\Delta(h)$ $=\sum_{(a)}(S(a_{(1)})\otimes 1)\Delta(a_{(2)}h)$
    $=\sum_{(a)}(S(a_{(1)})\varepsilon(a_{(2)})\otimes 1)\Delta(h)$
    $=(S(a)\otimes 1)\Delta(h).$
    The other equation is proven in a similar way.
In the case of a multiplier Hopf -algebra we can use that $S(a^{*})=S^{-1}(a)^{*}$ to obtain one equation from the other.
Since this is the first time where the Sweedler notation is used, let us explain it a bit more in the following remark.
1.6.
  • Remark If in all of the above expressions, we multiply in the first factor with any element from the left, we see that we have all these formulas within the algebraic tensor product of $A$ with itself. The various identities we use, like $m(\iota\otimes S)\Delta(a))=\varepsilon(a)1$ in the first equality, multiplication by such an element makes it possible to interpret these formulas in the correct way.
1.7.
  • Proposition Assume that $(A,\Delta)$ is a regular multiplier Hopf algebra. If $h$ is a left cointegral then the left and the right leg of $\Delta(h)$ are all of $A$.
  • Proof: Let $h$ be a left cointegral. Consider the subset $A_{0}$ of $A$ spanned by elements of the form $(\omega\otimes\iota)((c\otimes 1)\Delta(h))$ where $c\in A$ and $\omega$ in $A^{\prime}$. Assume that $\rho$ is a linear functional on $A$ that is 0 on this subset. Then $(\iota\otimes\rho)((c^{\prime}S(c)\otimes 1)\Delta(h))=0$ for all $c,c^{\prime}$. By the previous proposition we get $(\iota\otimes\rho)((c^{\prime}\otimes c)\Delta(h))=0$.
    Using the Sweedler notation we see that $\sum_{(h)}c^{\prime}h_{(1)}\rho(ch_{(2)})=0$. Then we apply the map $\Delta$ and obtain
    $\sum_{(h)}\Delta(c^{\prime})(h_{(1)}\otimes h_{(2)})\rho(ch_{(3)})=0.$
    We can multiply with an element $c^{\prime\prime}$ from the left in the first factor and use that $(A\otimes 1)\Delta(A)=A\otimes A$. Then we find
    $\sum_{(h)}ph_{(1)}\otimes qh_{(2)}\rho(ch_{(3)})=0$
    for all $p,q,c$. We apply the antipode on the second factor and multiply from the left with an element $r$. Then
    $\sum_{(h)}ph_{(1)}\otimes rS(h_{(2)})S(q)\rho(ch_{(3)})=0.$
    Now $h_{(1)}$ is covered by $p$ and $h_{(2)}$ is covered by $r$. Therefore we can cancel $S(q)$. We write this as
    $\sum_{(h)}ph_{(1)}\otimes rS(h_{(2)})\otimes\rho(\,\cdot\,h_{(3)})=0.$
    Finally we apply the evaluation map on the last two factors and we get for the left hand side
    $\sum_{(h)}ph_{(1)}\rho(rS(h_{(2)})h_{(3)})=\sum_{(h)}ph_{(1)}\varepsilon(h_{(2)})\rho(r)=ph\rho(r).$
    As this is now equal to 0 for all $p$ and $r$ we obtain that $\rho=0$. Because this is true for all linear functionals that are 0 on the subspace $A_{0}$ we must have $A_{0}=A$.
    ii) Next we claim that also the span of elements $(\omega(\,\cdot\,b)\otimes\iota)\Delta(h)$ is all of $A$. Indeed, if $\rho$ is 0 on all such elements, we must have $(\iota\otimes\rho)(\Delta(h)(b\otimes 1))=0$ for all $b$. We can multiply with $c$ from the left and cancel $b$ so that again $(\iota\otimes\rho)((c\otimes 1)\Delta(h))=0$ for all $c$. Then we can proceed as in item i).
    iii) In a similar way we can prove that the spaces spanned by elements of the form $(\iota\otimes\omega)((1\otimes b))\Delta(h)$ and $(\iota\otimes\omega)(\Delta(h)(1\otimes b))$ are both equal to all of $A$.
Also the two legs of $\Delta(k)$ are all of $A$ for a right cointegral $k$ in a regular multiplier Hopf algebra.
We can go a little further here, using a standard argument from the general theory of multiplier Hopf algebras.
1.8.
  • Proposition Let $(A,\Delta)$ be a regular multiplier Hopf algebra with a left cointegral $h$. Assume that $\omega$ is a faithful linear functional on $A$. Then any element of $A$ is of the form $(\omega(\,\cdot\,a)\otimes\iota)\Delta(h)$ for some $a\in A$.
  • Proof: Let $\rho$ be any linear functional on $A$ satisfying $\rho(c)=0$ for all elements $c$ of the form $c=(\omega(\,\cdot\,a)\otimes\iota)\Delta(h)$ for some $a\in A$. Then, for all $a,b$
    $(\omega(\,\cdot\,ab)\otimes\rho)(\Delta(h))=0$
    and so $\omega(xb)=0$ where $x=(\iota\otimes\rho)(\Delta(h)(a\otimes 1))$. Because this holds for all $b$ and as $\omega$ is faithful, it follows that $x=0$. Then $\omega_{1}(x)=0$ for all $\omega_{1}$. This means that $\rho$ is actually 0 on all elements of the form
    $(\omega_{1}(\,\cdot\,a)\otimes\iota)\Delta(h).$
    By Proposition 1.7 $A$ is spanned by elements of this form. Hence $\rho$ is 0 on all of $A$. This proves the result.
A similar result will be true with different variations, as well as for right cointegrals.
Here is an important consequence about the structure of the underlying algebra $A$.
1.9.
  • Proposition Let $(A,\Delta)$ be a regular multiplier Hopf algebra with a left cointegral $h$. Then $AaA$ is finite-dimensional for all $a$.
  • Proof: i) Consider $c\in A$ and write $\Delta(h)(c\otimes 1)=\sum_{i}p_{i}\otimes q_{i}$. For all $a$, we have $(1\otimes a)\Delta(h)=(S(a)\otimes 1)\Delta(h)$ by Proposition 1.5. Next let $\omega\in A^{\prime}$ and put $x=(\omega(\,\cdot\,c)\otimes\iota)\Delta(h)$. Then
    $ax$ $=(\omega(\,\cdot\,c)\otimes\iota)((1\otimes a)\Delta(h))$
    $=(\omega(\,\cdot\,c)\otimes\iota)((S(a)\otimes 1)\Delta(h))$
    $=\omega(S(a)p_{i})q_{i}$
    and we see that $ax$ belongs to the finite-dimensional space spanned by the elements $(q_{i})$. By Proposition 1.7 any element of $A$ is spanned by such elements $x$ and therefore we have shown that $aA$ is finite-dimensional for all $a$ in $A$.
    ii) We can apply a similar argument with a right cointegral and this gives that $Aa$ is finite-dimensional for all $a$ in $A$.
    iii) Combining the two results we can conclude that $AaA$ is finite-dimensional for all $a\in A$.
We can now prove the following characterization property.
1.10.
  • Theorem Assume that $(A,\Delta)$ is a multiplier Hopf -algebra of discrete type. If $A$ is an operator algebra then $A$ is a direct sum of finite-dimensional matrix algebras over $\mathbb{C}$.
  • Proof: For all $a$ we have a finite-dimensional two-sided ideal $AaA$. It contains $a$ because $A$ has local units. Because $A$ is an operator algebra, $AaA$ is the direct sum of full matrix algebras. Because any element $a$ belongs to such a sum of matrix algebras, we must have that $A$ is direct sum of matrix algebras.
The next section will be devoted to the study of multiplier Hopf -algebras of discrete type where the underlying -algebra is an operator algebra as in this theorem. They are what we call here discrete quantum groups.
Invariant integrals on $(A,\Delta)$
In what follows, we assume that $(A,\Delta)$ is a multiplier Hopf algebra of discrete type as in Definition 1.3. We also assume now that it is a regular multiplier Hopf algebra.
1.11.
  • Theorem Assume that $(A,\Delta)$ is a regular multiplier Hopf algebra with a left cointegral $h$. Then there exist non-zero linear functionals $\varphi$ and $\psi$ on $A$ satisfying
    $(\iota\otimes\varphi)\Delta(a)$ $=\varphi(a)1$ (1.1)
    $(\psi\otimes\iota)\Delta(a)$ $=\psi(a)1$ (1.2)
    for all $a\in A$. These functionals are faithful.
  • Proof: i) We will denote by $A^{\prime}_{r}$ the space of linear functionals on $A$ spanned by elements of the form $\omega(c\,\cdot\,)$ where $c\in A$ and $\omega\in A^{\prime}$. Define $\varphi(x)=\omega(1)$ if $x=(\omega\otimes\iota)\Delta(h)$ with $\omega\in A^{\prime}_{r}$. As we have seen in the proof of Proposition 1.7 above all elements in $A$ are of this form. Moreover, if for such an element we have $(\omega\otimes\iota)\Delta(h)=0$ we must have $\omega=0$ because also the left leg of $\Delta(h)$ is all of $A$.
    Now take any $a\in A$ and write $a=(\omega\otimes\iota)\Delta(h)$ with $\omega\in A^{\prime}_{r}$. Then for all $\omega^{\prime}\in A^{\prime}_{r}$ we get
    $\varphi((\omega^{\prime}\otimes\iota)\Delta(a))$ $=\varphi((\omega^{\prime}\otimes\iota)(\omega\otimes 1\otimes 1)(\Delta\otimes\iota)\Delta(h))$
    $=\varphi((((\omega\otimes\omega^{\prime})\circ\Delta)\otimes\iota)\Delta(h))$
    $=(\omega\otimes\omega^{\prime})\Delta(1)=\omega(1)\omega^{\prime}(1)$
    $=\omega^{\prime}(1)\varphi(a).$
    Hence $(\iota\otimes\varphi)\Delta(a)=\varphi(a)1$. We also have $\varphi(h)=1$ by definition.
    ii) The case of the right integral $\psi$ is similar. If we use a right cointegral $k$ to construct it, we get $\psi(k)=1$.
    iii) Non-zero invariant linear functionals on a regular multiplier Hopf algebra are automatically faithful.
We have used coassociativity of $\Delta$ and the extension of $\Delta$ to the multiplier algebra $M(A)$. All of this is justified because $\omega,\omega^{\prime}\in A^{\prime}_{r}$.
This result, as well as some of the following results, are also found in [VD-Zh0].
1.12.
  • Proposition Let $(A,\Delta)$ be a regular multiplier Hopf algebra of discrete type. Let $h$ be a left cointegral and put $k=S^{-1}(h)$. Let $\varphi$ be the left integral satisfying $\varphi(h)=1$ and $\psi$ the right integral satisfying $\psi(k)=1$. Then $\varphi\circ S=\psi$. If $\varepsilon(h)=1$ we also have $\psi\circ S=\varphi$.
  • Proof: i) By assumption $S(k)=h$ and $\varphi(S(k))=\varphi(h)=1$. Because $\varphi\circ S$ is a right integral and takes the same value in $k$ as $\psi$, we will have $\psi=\varphi\circ S$ by the uniqueness of integrals.
    ii) By the uniqueness of cointegrals, we have that $S^{2}(h)$ is a scalar multiple of $h$. Now $\varepsilon(S^{2}(h))=\varepsilon(h)$. Therefore, if $\varepsilon(h)$ is non-zero, we must have $S^{2}(h)=h$. Then also S(h)=k and $\varphi=\psi\circ S$ using the argument as in i).
We see that the scaling constant $\nu$, defined by $\varphi\circ S^{2}=\nu\varphi$, has to be equal to $1$ if the cointegral $h$ satisfies $\varepsilon(h)\neq 0$.
If $(A,\Delta)$ is multiplier Hopf -algebra of discrete type and $A$ is an operator algebra, we will show in the next section that the integrals are positive.
The modular automorphisms
From the general theory of algebraic quantum groups (i.e. regular multiplier Hopf algebras with integrals) we know the existence of an automorphism $\sigma$ of $A$ such that $\varphi(ab)=\varphi(b\sigma(a))$ for all $a,b\in A$ when $\varphi$ is a left integral. For the integrals we have obtained in Theorem 1.11 we can obtain the following property of the modular automorphisms.
1.13.
  • Proposition Let $h$ be a left cointegral and assume that it is also a right cointegral. Then for all $a,b\in A$ we have
    $\varphi(ab)=\varphi(bS^{2}(a))\qquad\text{and}\qquad\psi(ab)=\psi(bS^{-2}(a)).$
  • Proof: i) Because $h$ is a left cointegral, we have $(1\otimes a)\Delta(h)=(S(a)\otimes 1)\Delta(h)$ for all $a$ and because we assume that it is also a right cointegral, we also have $\Delta(h)(a\otimes 1)=\Delta(h)(1\otimes S(a))$ for all $a$.
    ii) Now let $a\in A$. Then
    $(\iota\otimes\varphi)((1\otimes a)\Delta(h))$ $=(\iota\otimes\varphi)((S(a)\otimes 1)\Delta(h))$
    $=S(a)=(\iota\otimes\varphi)(\Delta(h)(S(a)\otimes 1)$
    $=(\iota\otimes\varphi)(\Delta(h)(1\otimes S^{2}(a))).$
    Because the right leg of $\Delta(h)$ is all of $A$ we get $\varphi(ab)=\varphi(bS^{2}(a))$ for all $b\in A$.
    iii) Similarly we get $\psi(ab)=\psi(bS^{-2}(a))$.
1.14.
  • Proposition If $h$ is a left cointegral and also a right cointegral we have $\sigma(h)=h$. Then also $S^{2}(h)=h$.
  • Proof: For all $a\in A$ we have $\varphi(ha)=\varepsilon(a)\varphi(h)=\varphi(ah)$ and therefore $\sigma(h)=h$. Because $S^{2}$ and $\sigma$ coincide we also get $S^{2}(h)=h$.
The property $S^{2}(h)=h$ is true for any left cointegral if $\varepsilon(h)\neq 0$. In that case, $h$ is also a right cointegral. What we obtain here is that $S^{2}(h)=h$ holds if $h$ is also a right cointegral. It is not clear if this would imply already that $\varepsilon(h)\neq 0$.
We also have a converse result.
1.15.
  • Proposition Assume that $(A,\Delta)$ is a regular multiplier Hopf algebra with a left cointegral $h$. Let $\sigma$ be the modular automorphism of the left integral $\varphi$. If $\sigma(h)=h$ we must have that $h$ is also a right cointegral.
  • Proof: If $\sigma(h)=h$ we have $\varphi(ha)=\varphi(ah)=\varepsilon(a)\varphi(h)=\varepsilon(a)$. This means that
    $\varphi(hab)=\varepsilon(ab)=\varepsilon(a)\varepsilon(b)=\varepsilon(a)\varphi(hb).$
    From the faithfulness of $\varphi$ we find that $ha=\varepsilon(a)h$.
The modular element
Again from the general theory, we know the existence of a multiplier $\delta\in M(A)$ given by $\psi(a)=\varphi(a\delta)$ for all $a$ (assuming that the $\psi$ and $\varphi$ are normalized in such a way that $\psi=\varphi\circ S$). We also have $(\varphi\otimes\iota)\Delta(a)=\varphi(a)\delta$ for all $a$. In particular, because we use the normalization $\varphi(h)=1$, we get here $\delta=(\varphi\otimes\iota)\Delta(h)$. Again we can not say much more except when, as in Proposition 1.15 , a left cointegral is also a right cointegral.
1.16.
  • Proposition If $h$ is also a right cointegral, then for all $a$ we have $S^{4}(a)=\delta^{-1}a\delta$.
  • Proof: For $a\in A$ we have
    $\delta a$ $=(\varphi\otimes\iota)((1\otimes a)\Delta(h))$
    $=(\varphi\otimes\iota)((S(a)\otimes 1)\Delta(h))$
    $=(\varphi\otimes\iota)(\Delta(h)(S^{3}(a)\otimes 1)$
    $=(\varphi\otimes\iota)(\Delta(h)(1\otimes S^{4}(a))$
    $=\delta S^{4}(a).$
This is Radford’s formula proven in [Ra0] for the fourth power of the antipode because $\widehat{\delta}=1$. That $\widehat{\delta}=1$ follows from the general theory. Indeed, we know that $\langle a,\widehat{\delta}\rangle=\varepsilon(\sigma^{-1}(a))$, see e.g. Proposition 2.3 in [VD-part1] or Proposition 1.2.3 in [VD-lnalg]. In this case we have $S^{2}=\sigma$ and because $\varepsilon(S^{-2}(a))=\varepsilon(a)$ we get $\langle a,\widehat{\delta}\rangle=\varepsilon(a)$ and therefore $\widehat{\delta}=1$.
The element $\Delta(h)$ as a separability idempotent
In this item, we again assume that $h$ is a left cointegral and that it satisfies $\varepsilon(h)=1$. Then it also a right cointegral and we can assume that it is an idempotent as we saw in Proposition 1.2. It follows that $\Delta(h)$ is an idempotent in $M(A\otimes A)$. It plays and important role in the next section on discrete quantum groups.
1.17.
  • Proposition The element $\Delta(h)$ is a (regular) separability idempotent in $M(A\otimes A)$ in the sense of [VD-si1].
  • Proof: i) First we look at the setting in [VD-si1]. For the algebras $B$ and $C$ we take $A$ and we let $E=\Delta(h)$. The algebra $A$ is non-degenerate and $\Delta(h)$ is an idempotent in $M(A\otimes A)$. Because $A$ is a regular multiplier Hopf algebra we know that elements of the form $\Delta(h)(1\otimes a)$ and $\Delta(h)(a\otimes 1)$ belong to $A\otimes A$ for all $a\in A$ and similarly with $a$ on the other side. This shows that Assumption 1.1 of [VD-si1] is fulfilled and that $\Delta(h)$ is regular in the sense of that assumption.
    ii) We have shown that the legs of $\Delta(h)$ are all of $A$ and hence it is full. Also Proposition 1.4 in [VD-si1] here says that $\Delta(h)(1\otimes a)=0$ only if $a=0$, and similarly $(a\otimes 1)\Delta(h)=0$ only if $a=0$. This we knew already because we have a multiplier Hopf algebra and $h\neq 0$.
    iii) Finally we come to Definition 1.5 of [VD-si1]. Since we know here already that
    $\Delta(h)(a\otimes 1)=\Delta(h(1\otimes S(a))\qquad\quad\text{and}\qquad\quad(1\otimes a)\Delta(h)=(S(a)\otimes 1)\Delta(h)$
    for all $a$, and because $S$ is bijective, we finally get that $\Delta(h)$ satisfies all the requirements and so it is a regular separability idempotent.
The maps $S:B\to C$ and $S^{\prime}:C\to B$ we have in Proposition 1.7 of [VD-si1] are precisely the antipode $S$ of $A$. If we look at the Proposition 1.8 in [VD-si1], we get here
$m(S\otimes\iota)(\Delta(h)(1\otimes a))=\sum_{(h)}S(h_{(1)})h_{(2)}a=\varepsilon(h)a$
and consequently $m(S\otimes\iota)(\Delta(h)(1\otimes a))=a$ for all $a$.
Similarly
$m(\iota\otimes S)((a\otimes 1)\Delta(h))=\sum_{(h)}ah_{(1)})S(h_{(2)})=\varepsilon(h)a$
and consequently $m(\iota\otimes S)((\otimes\iota)\Delta(h))=a$ for all $a$.
1.18.
  • Remark In the case where the left cointegral $h$ satisfies $\varepsilon(h)=1$, so that $\Delta(h)$ is a separability idempotent, we can apply general results from [VD-si1] to obtain the existence of $\varphi$ and $\psi$, as in Theorem 1.11 above. However, we only would get $(\iota\otimes\varphi)\Delta(h)=1$ and $(\psi\otimes\iota)\Delta(h)=1$ and we would still need the argument given in the proof of Theorem 1.11 to obtain left and right invariance on all of $A$.
    Also the results in Proposition 1.13 and Proposition 1.16 already follow from the general theory in [VD-si1].
In the next section, we will give an explicit formula for $\Delta(h)$ and refer to this item on separability idempontents.
2 . Discrete quantum groups
In this paper, a discrete quantum group is viewed as a discrete quantum space with a comultiplication. Therefore we first give the definition of a discrete quantum space as we view it here.
Discrete quantum spaces
Let $I$ be an index set. Associate to each member $\alpha$ of $I$ a natural number $n(\alpha)$ and the -algebra $A_{\alpha}$ of $n(\alpha)\times n(\alpha)$ complex matrices (with the usual -algebra structure). Let $A$ be the (algebraic) direct sum of these algebras. Elements of $A$ are written as $(a_{\alpha})_{\alpha\in I}$ where $a_{\alpha}\in A_{\alpha}$ for each $\alpha\in I$. By assumption only finitely many components $a_{\alpha}$ are non-zero. This direct sum $A$ is again a -algebra in the obvious sense. Observe that the algebra has no unit, except when $I$ is a finite set, but that the product is non-degenerate (as a bilinear form). We say that the algebra $A$ is a direct sum of matrix algebras. The algebras $A_{\alpha}$ are considered as sitting inside $A$. The common notation is $\sum_{\alpha\in I}\oplus A_{\alpha}$.
Elements in the multiplier algebra $M(A)$ of $A$ can be written as $(a_{\alpha})_{\alpha\in I}$ where $a_{\alpha}\in A_{\alpha}$ for each $\alpha\in I$, but without further restrictions. This algebra is called the direct product of the algebras $A_{\alpha}$. Here the usual notation is $\prod_{\alpha\in I}A_{\alpha}$.
We also consider the tensor product $A\otimes A$ of $A$ with itself and the multiplier algebra $M(A\otimes A)$ of $A\otimes A$. The algebra $A\otimes A$ is the direct sum of the tensor product algebras $A_{\alpha}\otimes A_{\beta}$ where $\alpha,\beta\in I$ and of course the multiplier algebra $M(A\otimes A)$ is the direct product of all these algebras. So elements in $M(A\otimes A)$ are written as $(a_{\alpha\beta})_{\alpha,\beta\in I}$ with $a_{\alpha\beta}\in A_{\alpha}\otimes A_{\beta}$. When only finitely many components are non-zero, we get the elements of $A\otimes A$.
The identity in $A_{\alpha}$ will be denoted by $1_{\alpha}$ and it will also be considered as a minimal central projection in $A$.
Now we come to quantizing the ’product structure’. So we introduce the notion of a coproduct on such an algebra.
Discrete quantum groups
2.1.
  • Definition Let $A$ be a direct sum of matrix algebras as above. A coproduct (or comultiplication) on $A$ is a -homomorphism $\Delta:A\to M(A\otimes A)$ such that
    i) $\Delta(A)(1\otimes A)\subseteq A\otimes A$ and $(A\otimes 1)\Delta(A)\subseteq A\otimes A$,
    ii) $(\iota\otimes\Delta)\Delta(a)=(\Delta\otimes\iota)\Delta(a)$ for all $a\in A$.
Recall that we use $\iota$ to denote the identity map.
The condition i) is used to give a meaning to the condition in ii). Indeed, if we multiply the left hand side of the equation on the left with an element of the form $c\otimes 1\otimes 1$ where $c\in A$, the expression is $(\iota\otimes\Delta)((c\otimes 1)\Delta(a))$ and this is well-defined in $M(A\otimes A\otimes A)$. Similarly, we multiply the right hand side of the equation from the right with an element of the form $1\otimes 1\otimes b$ with $b\in A$ and then the expression is $(\Delta\otimes\iota)(\Delta(a)(1\otimes b))$. This is how the assumptions in i) are used to interpret the equation in ii) in the multiplier algebra $M(A\otimes A\otimes A)$.
Condition ii) is called coassociativity.
The above definition of a coproduct is also possible when $A$ is any -algebra with a non-degenerate product (see [VD-mha]). If we do not assume the conditions in i), but only assume that $\Delta$ is non-degenerate (i.e. that $\Delta(A)(A\otimes A)=A\otimes A$), it is still possible to give a meaning to coassociativity, see e.g. the note on coassociativity [VD-coass, VD-coass1]. This is not so hard to see in the case of a direct sum. However, as we will see, the conditions in i) are quite natural and similar conditions are often taken as part of the notion of a coproduct itself.
Then, the following is the main definition for our approach here.
$2.$
  • Definition A discrete quantum group is a pair $(A,\Delta)$ of a -algebra AA that is a direct sum of matrix algebras and a coproduct $\Delta$ on A (as in Definition 2.1) such that $(A,\Delta)$ is a multiplier Hopf -algebra.
We know that there is a unique counit and a unique antipode $S$. The counit is defined as a linear map $\varepsilon:A\to\mathbb{C}$ satisfying
$(\varepsilon\otimes\iota)\Delta(a)=a\qquad\text{and}\qquad(\iota\otimes\varepsilon)\Delta(a)=a$
for all $a$ in $A$. The antipode is defined as a linear map $S:A\to A$ such that
$m(S\otimes\iota)\Delta(a)=\varepsilon(a)1\qquad\text{and}\qquad m(\iota\otimes S)\Delta(a)=\varepsilon(a)1$
for all $a$ in $A$ (where $m$ is the multiplication map, seen as a linear map from $A\otimes A$ to $A$).
Also these formulas make sense in the multiplier algebra $M(A)$. One just has to multiply, left or right (depending on the case) with an element in $A$.
From the fact that the comultiplication is a homomorphism, it follows that $\varepsilon$ is a homomorphism and that $S$ is a anti-homomorphism. Because $\Delta$ is a -homomorphism, it follows that $\varepsilon$ is also a -homomorphism and that $S$ satisfies the equation $S(S(a)^{*})^{*}=a$ for all $a$ in $A$. In particular, $S$ is bijective and $S(a^{*})=S^{-1}(a)^{*}$ for all $a$. We see that $S$ is a -map if and only if $S^{2}=\iota$.
Before we continue, we give the basic example, motivating the main definition. It shows that the counit is indeed the dual of the unit element in a group and that the antipode is what corresponds to the inverse in a group. The example justifies Definition 2.2 as a natural definition of a discrete quantum group.
2.3.
  • Example Let $G$ be a (discrete) group. Consider the algebra of complex functions on $G$ with finite support (with pointwise operations). Define a coproduct $\Delta$ on $A$ by $\Delta(f)(p,q)=f(pq)$ where $p,q\in G$. Then $(A,\Delta)$ is a discrete quantum group.
In this case, the index set is equal to $G$ and $n(p)=1$ so that $A_{p}=\mathbb{C}1$ for all $p\in G$.
In Section 4 we also consider some other important special cases and we refer to [VD-suq2] for a more specific example.
There is a lot more to say about this definition. This is done in the appendix. But we give here some comments already.
2.4.
  • Remark i) If $(A,\Delta)$ is a pair of a -algebra with a coproduct as in Definition 2.2, we can look at the linear maps $T_{1}$ and $T_{2}$ defined on $A\otimes A$ by
    $T_{1}(a\otimes b)=\Delta(a)(1\otimes b)\qquad\text{and}\qquad T_{2}(a\otimes b)=(a\otimes 1)\Delta(b).$
    By the assumption on the coproduct, they map into $A\otimes A$. It can be shown however that they are bijective when a counit and an antipode exists. In that case, the inverses are given by
    $T_{1}^{-1}(a\otimes b)$ $=((\iota\otimes S)\Delta(a))(1\otimes b)$ (2.1)
    $T_{2}^{-1}(a\otimes b)$ $=(a\otimes 1)((S\otimes\iota)\Delta(b)).$ (2.2)
    So, the pair $(A,\Delta)$ is a multiplier Hopf -algebra as defined in [VD-mha].
    ii) Conversely, if $(A,\Delta)$ is a multiplier Hopf -algebra and if $A$ is a direct sum of matrix algebras, then it is a discrete quantum group in the sense of Definition 2.2. The reason is that in any multiplier Hopf -algebra, there exists a counit and an antipode (see [VD-mha].
So, a discrete quantum group is nothing else but a multiplier Hopf -algebra where the underlying algebra is a direct sum of matrix algebras (cf. the definition given of a discrete quantum group in [VD-discrete]).
In the previous section, we developed the theory of multiplier Hopf algebras of discrete type. In Theorem 1.10, we have conditions to assure that the underlying algebra is a direct sum of matrix algebras. In other words, we get a discrete quantum group (as in definition 2.1). Next we will see that also conversely, if we have a discrete quantum group, it has cointegrals and so it is a multiplier Hopf algebra of discrete type (as in Definition 1.3.
The cointegral in a discrete quantum group
The following result is fairly obvious and we only include a proof for completeness.
2.5.
  • Proposition Assume that $A$ is a direct sum of matrix algebras as considered before and let $\varepsilon$ be a -homomorphism from $A$ to $\mathbb{C}$. Then there exists a unique non-zero self-adjoint idempotent $h$ in $A$ with the property that $ah=\varepsilon(a)h$ for all $a\in A$. We also have $ha=\varepsilon(a)h$ for all $a\in A$.
  • Proof: The kernel $A_{0}$ of $\varepsilon$ is a two-sided, -invariant ideal of $A$ with codimension $1$. As $A$ is a direct sum of matrix algebras, we must have a central projection $p\in M(A)$ such that $A_{0}=Ap$. When h=1ph=1-p, then $h\in A$ and $\varepsilon(h)=1$ and so $\varepsilon(ah-\varepsilon(a)h)=0$. This means that $ah-\varepsilon(a)h$ belongs to the kernel $Ap$ for all $a$. However, because $hp=0$, also $ah-\varepsilon(a)h=0$ and so $ah=\varepsilon(a)h$. By taking adjoints, we get also $ha=\varepsilon(a)h$ for all $a$. If now $k$ is another such element, then $hk=\varepsilon(h)k=k$ and also $hk=\varepsilon(k)h=h$ and therefore $h=k$.
Remark that in the proof above, we have applied $\varepsilon$ on $M(A)$. This is no problem as we can easily extend $\varepsilon$ to a -homomorphism on $M(A)$ by $\varepsilon(m)=\varepsilon(am)$ where $a$ is in $A$ and satisfies $\varepsilon(a)=1$.
If $(A,\Delta)$ is a discrete quantum group as in Definition 2.2, we not only have that $h^{*}=h$ and $h^{2}=h$, but also $S(h)=h$ because $\varepsilon(S(a))=\varepsilon(a)$ for all $a$.
A trivial consequence is the following.
2.6.
  • Proposition When $(A,\Delta)$ is a discrete quantum group, then it is a multiplier Hopf algebra of discrete type. The left contegral $h$ satisfies $\varepsilon(h)=1$, it is an idempotent and $S(h)=h$. It is also a right cointegral.
It turns out to give the integral on the dual - see Proposition 3.6 in the next section.
We can apply all the results of the previous section because $\varepsilon(h)\neq 0$.
2.7.
  • Notation i) The support of the counit $\varepsilon$ is a distinct one-dimensional component. We will denote the corresponding index with $e$, reminding us of the identity element in a group.
    ii) Let $\alpha$ be any index and consider the corresponding component $A_{\alpha}$. Because $S$ is an anti-isomorphism we have an index $\overline{\alpha}$ such that $S(A_{\alpha})=A_{\overline{\alpha}}$.
For $a\in A_{\alpha}$ we have $S(S(a_{\alpha})^{*})=a^{*}$. Because all the components are -invariant, this implies $S(A_{\overline{\alpha}})=A_{\alpha}$. In particular, $S^{2}(A_{\alpha})=A_{\alpha}$ for all $\alpha$. We will study these maps later in greater detail.
First we have the following result about the element $\Delta(h)$ where $h$ is the cointegral given in Proposition 2.5. Recall that we use $1_{\alpha}$ for the identity of $A_{\alpha}$ and consider it as a central element in $A$.
2.8.
  • Proposition We have
    $\Delta(h)(1\otimes 1_{\alpha})=\Delta(h)(1_{\overline{\alpha}}\otimes 1)=\Delta(h)(1_{\overline{\alpha}}\otimes 1_{\alpha}).$
This follows from the property that $S(1_{\alpha})=1_{\overline{\alpha}}$ for all $\alpha$.
The maps $a\mapsto S(a)^{*}$ and $a\mapsto S^{2}(a)$
There are plenty of traces on the algebra $A$, but on each component, the trace is unique up to a scalar. We will further work with the standard traces. Recall that $A_{\alpha}$ is the algebra of $n(\alpha)\times n(\alpha)$ matrices over the complex numbers.
2.9.
  • Notation Denote by $\omega_{\alpha}$ the unique trace on $A_{\alpha}$ satisfying $\omega_{\alpha}(1_{\alpha})=n(\alpha)$. We use $\omega$ for the trace on $A$ with these components on $A_{\alpha}$.
2.10.
  • Notation Denote $q=(\omega\otimes\iota)\Delta(h)$ and $q_{\alpha}$ for the components of $q$. Because of Proposition 2.8 we have $q_{\alpha}=(\omega_{\overline{\alpha}}\otimes 1)\Delta(h)$.
The element $q$ is defined in the multiplier algebra $M(A)$.
2.11.
  • Proposition The element $q$ is positive and invertible in $M(A)$. Further $aq=qS^{2}(a)$ for all $a\in A$.
  • Proof: i) Using the formulas in Proposition 1.5 and the fact that $\omega$ is a trace, we find
    $aq$ $=(\omega\otimes\iota)((1\otimes a)\Delta(h))$
    $=(\omega\otimes\iota)((S(a)\otimes 1)\Delta(h))$
    $=(\omega\otimes\iota)(\Delta(h)(S(a)\otimes 1))$
    $=(\omega\otimes\iota)(\Delta(h)(1\otimes S^{2}(a)))$
    $=qS^{2}(a)$
    for all $a$.
    ii) For $a\in A_{\alpha}$ we get $aq_{\alpha}=q_{\alpha}S^{2}(a)$. Because $q_{\alpha}\neq 0$ we have that it is invertible. Because $\Delta(h)$ is positive and the trace is positive, we have that $q$ is positive.
In a similar way we can consider $(\iota\otimes\omega)\Delta(h)$. It will be a positive invertible element $p\in M(A)$ satisfying $ap=S^{-2}(a)p$ for all $a\in a$.
It is natural for $S^{2}$ to be implemented by a self-adjoint invertible element because of the fact that $S^{2}(a)^{*}=S^{-2}(a^{*})$ for all $a$, but that it can be implemented by a positive invertible elements is not automatic.
We will see later that $\omega_{\alpha}(q_{\alpha})=1$.
The scaling group and the unitary antipode
It has been shown in [Ku] that algebraic quantum groups admit an analytic structure. In particular, there is a one-parameter group $(\tau_{t})$ of automorphisms of $A$ with the property that all elements in $A$ are analytic. It is characterized by the fact that $S^{2}=\tau_{-i}$. This is called the scaling group. There is also a unitary antipode $R$ such that $S=R\tau_{-\frac{i}{2}}$. This is thought of as the polar decomposition of the antipode. These results are also found in [Ku-VD] and in the more recent survey paper [VD-part3]. The polar decomposition of the antipode is also a feature of the more general locally compact groups, see the references given in the corresponding item in the introduction.
In the case of a discrete quantum group, we get the following result.
2.12.
  • Proposition We get $\tau_{t}(a)=q^{-it}aq^{it}$ for all $a\in A$.
  • Proof: From $S=R\tau_{-\frac{i}{2}}$ we get $S^{2}=\tau_{-i}$. We use that $R^{2}=\iota$ and that $R$ and $\tau_{t}$ commute. Then, because $S^{2}(a)=q^{-1}aq$, we must have $\tau_{t}(a)=q^{-it}aq^{it}$.
The scaling group leaves all components of $A$ globally invariant and $\tau_{t}(a)=q_{\alpha}^{-it}aq_{\alpha}^{it}$ for all $a\in A_{\alpha}$.
The situation is different for the unitary antipode $R$. For this we get the following results. Recall that for each index $\alpha$ there is an index $\overline{\alpha}$ such that $S:A_{\alpha}\to A_{\overline{\alpha}}$.
2.13.
  • Lemma Take any vector $\xi_{0}$ in $\mathcal{H}_{\alpha}\otimes\mathcal{H}_{\overline{\alpha}}$. Define a conjugate linear map $x:\mathcal{H}_{\alpha}\to\mathcal{H}_{\overline{\alpha}}$ by
    $\langle x\xi,\eta\rangle=\langle\Delta(h)\xi_{0},\xi\otimes\eta\rangle.$ (2.3)
    Then $xa=S(a)^{*}x$ and $x^{*}xa=S^{2}(a)x^{*}x$ for all $a\in A_{\alpha}$.
  • Proof: i) Recall that $A_{\alpha}$ acts on the finite-dimensional space $\mathcal{H}_{\alpha}$ while $A_{\overline{\alpha}}$ acts on $\mathcal{H}_{\overline{\alpha}}$. Then $\Delta(h)\xi_{0}$ again belongs to $\mathcal{H}_{\alpha}\otimes\mathcal{H}_{\overline{\alpha}}$. And because these spaces are finite-dimensional, the formula (2.3) defines a conjugate linear map from $\mathcal{H}_{\alpha}$ to $\mathcal{H}_{\overline{\alpha}}.$
    ii) Now for $a\in A_{\alpha}$ we get
    $\langle xa\xi,\eta\rangle$ $=\langle\Delta(h)\xi_{0},a\xi\otimes\eta\rangle$
    $=\langle(a^{*}\otimes 1)\Delta(h)\xi_{0},\xi\otimes\eta\rangle$
    $=\langle(1\otimes S^{-1}(a^{*}))\Delta(h)\xi_{0},\xi\otimes\eta\rangle$
    $=\langle(1\otimes S(a)^{*})\Delta(h)\xi_{0},\xi\otimes\eta\rangle$
    $=\langle\Delta(h)\xi_{0},\xi\otimes S(a)\eta\rangle$
    $=\langle x\xi,S(a)\eta\rangle=\langle S(a)^{*}x\xi,\eta\rangle$
    and we see that $xa=S(a)^{*}x$.
    iii) If we take adjoints we get $x^{*}S(a)=a^{*}x^{*}$. With $b=S(a)$ we get
    $x^{*}b=S^{-1}(b)^{*}=S(b^{*}).$
    Now with $b=S(c)^{*}$ we find $x^{*}S(c)^{*}=S(S(c))x^{*}$. Combining this with the original formula we get
    $x^{*}xa=x^{*}S(a)^{*}x=S^{2}(a)x^{*}x.$
2.14.
  • Proposition There exists a vector $\xi_{0}$ in $\mathcal{H}_{\alpha}\otimes\mathcal{H}_{\overline{\alpha}}$ such that for the corresponding operator $x$ we have $x^{*}x=q_{\alpha}^{-1}$.
  • Proof: We have seen before that $aq_{\alpha}=q_{\alpha}S^{2}(a)$. It follows that $x^{*}x$ must be a scalar multiple of $q_{\alpha}^{-1}$. Because $\Delta(h)$ is non-zero, we have elements $\xi_{0}$ such that $\Delta(h)\xi_{0}\neq 0$ and $x\neq 0$. Therefore there exists a vector $\xi_{0}$ such that actually $x^{*}x=q_{\alpha}^{-1}$.
Consider the polar decomposition of $x$. It is of the form $x=J_{\alpha}q_{\alpha}^{-\frac{1}{2}}$ where $J_{\alpha}$ is a conjugate linear isomorphism of $\mathcal{H}_{\alpha}$ to $\mathcal{H}_{\overline{\alpha}}$.
For the unitary antipode $R$ we have the following.
2.15.
  • Proposition The unitary anitpode $R$ maps $A_{\alpha}$ to $A_{\overline{\alpha}}$ and for $a\in A_{\alpha}$ we have $R(a)=J_{\alpha}(a^{*})J_{\alpha}$.
  • Proof: For $a\in A_{\alpha}$ we have
    $S(a)^{*}=xax^{-1}=J_{\alpha}q_{\alpha}^{-\frac{1}{2}}aq_{\alpha}^{\frac{1}{2}}J_{\alpha}.$
    so that $S(a)=J_{\alpha}(q_{\alpha}^{-\frac{1}{2}}aq_{\alpha}^{\frac{1}{2}})^{*}J_{\alpha}$. On the other hand, by definition
    $S(a)=R(\tau_{-\frac{i}{2}}(a))=R(q_{\alpha}^{-\frac{1}{2}}aq_{\alpha}^{\frac{1}{2}}).$
    It follows that $R(c)=J_{\alpha}c^{*}J_{\alpha}$ for all $c\in A_{\alpha}$.
A formula for the components of $\Delta(h)$
In this item we fix an index $\alpha$ in $I$.
Now we choose matrix elements $(e_{ij})$ in $A_{\alpha}$ and we assume that $q_{\alpha}=\sum_{i}\lambda_{i}e_{ii}$. This is possible because $q_{\alpha}$ is a positive element. All the eigenvalues $\lambda_{i}$ are strictly positive as $q_{\alpha}$ is invertible.
2.16.
  • Proposition We have
    $\Delta(h)(1_{\overline{\alpha}}\otimes 1_{\alpha})=\sum_{ij}\lambda_{i}\,S(e_{ij})\otimes e_{ji}.$
  • Proof: Define $H$ in $A_{\overline{\alpha}}\otimes A_{\alpha}$ by
    $H=\sum_{ij}\lambda_{i}\,S(e_{ij})\otimes e_{ji}.$
    i) First we show that $H(a\otimes 1)=H(1\otimes S(a))$ for all $a$. For any matrix element $e_{k\ell}$ we find
    $H(1\otimes e_{k\ell})$ $=\sum_{i,j}\lambda_{i}\,S(e_{ij})\otimes e_{ji}e_{k\ell}$
    $=\sum_{j}\lambda_{k}S(e_{kj})\otimes e_{j\ell}$
    $H(S^{-1}(e_{k\ell})\otimes 1)$ $=\sum_{i,j}\lambda_{i}\,S(e_{ij})S^{-1}(e_{k\ell})\otimes e_{ji}$
    $=\sum_{i,j}\lambda_{i}\lambda_{\ell}^{-1}\lambda_{k}S(e_{ij})S(e_{k\ell})\otimes e_{ji}$
    $=\sum_{i,j}\lambda_{i}\lambda_{\ell}^{-1}\lambda_{k}S(e_{k\ell}e_{ij})\otimes e_{ji}$
    $=\sum_{j}\lambda_{\ell}\lambda_{\ell}^{-1}\lambda_{k}S(e_{kj})\otimes e_{j\ell}$
    $=\sum_{j}\lambda_{k}S(e_{kj})\otimes e_{j\ell}.$
    We have used that $qa=S^{2}(a)q$ and that $q=\sum_{i}\lambda_{i}e_{ii}$. We see that indeed
    $H(1\otimes e_{k\ell})=H(S^{-1}(e_{k\ell})\otimes 1).$
    Then $H(S^{-1}(a)\otimes 1)=H(1\otimes a)$ for all $a\in A_{\alpha}$ and so also $H(a\otimes 1)=H(1\otimes S(a))$ for all $a\in A_{\overline{\alpha}}$.
    ii) Then we get
    $H\Delta(h)$ $=\sum_{(h)}H(h_{(1)}\otimes h_{(2)})$
    $=\sum_{(h)}H(1\otimes S(h_{(1)})h_{(2)})$
    $=\varepsilon(h)H=H.$
    On the other hand
    $H\Delta(h)$ $=\sum_{i,j}\lambda_{i}(S(e_{ij})\otimes e_{ji})\Delta(h)$
    $=\sum_{i,j}\lambda_{i}(1\otimes e_{ji}e_{ij})\Delta(h)$
    $=\sum_{i,j}\lambda_{i}(1\otimes e_{jj})\Delta(h)$
    $=c\Delta(h)(1\otimes 1_{\alpha}).$
    where $c=\sum_{i}\lambda_{i}$.
    iii) It follows that $H=c\Delta(h)(1_{\overline{\alpha}}\otimes 1_{\alpha})$. It is only natural that this constant is present because we have only used that $q$ implements $S^{2}$ and that is then true for any scalar multiple of $q$. However, we have a specific element $q$ in $A_{\alpha}$, defined as $(\omega\otimes\iota)\Delta(h)$. If we apply the standard trace $\omega$ on the first factor on the left we have
    $(\omega\otimes 1)H=\sum_{ij}\lambda_{i}\,\omega(S(e_{ij}))\otimes e_{ji}=\sum_{i}\lambda_{i}e_{ii}=q.$
    On the right we find $cq$ so that $c=1$ and hence $H=\Delta(h)(1_{\overline{\alpha}}\otimes 1_{\alpha})$. This completes the proof.
The number $c$ in the proof above is nothing else but $\omega_{\alpha}(q)$ and so this is $1$ as we announced already. It gives in fact that every component of $\Delta(h)$ in $A_{\overline{\alpha}}\otimes A_{\alpha}$ is a one-dimensional projection. This result is obtained in the original paper, see Proposition 4.4 in [VD-discrete].
We are now ready for the main property here. We refer to our papers on separability idempotents [VD-si1, VD-si2].
2.17.
  • Proposition The element $\Delta(h)$ is a separability idempotent in $M(A\otimes A)$.
  • Proof: i) We first look at the setting in [VD-si1]. For both the algebras $B$ and $C$ in that paper, we simply take $A$, and for $E$ we take $\Delta(h)$. The algebra $A$ is non-degenerate and $\Delta(h)$ is an idempotent in $M(A\otimes A)$. Here we have -algebras and $\Delta(h)$ is self-adjoint. We know that $\Delta(h)(1\otimes a)$ and $(a\otimes 1)\Delta(h)$ both belong to $A\otimes A$ and similarly on the other side. This shows that Assumption 1.1 of [VD-si1] are fulfilled. In particular $\Delta(h)$ is regular in the sense of that assumption.
    ii) We claim that $\Delta(h)$ is full in the sense that the legs of $\Delta(h)$ are all of $A$. This is proven in the paper, but follows more easily from the equation
    $H=\sum_{i}\lambda_{i}\,\Delta(h)(1_{\overline{\alpha}}\otimes 1_{\alpha})$
    Indeed, if $V$ is a subspace of $A$ satisfying $\Delta(h)(1\otimes a)\subseteq V\otimes A$ we get, with $a=e_{pq}$ from
    $\sum_{ij}\lambda_{i}\,S(e_{ij})\otimes e_{ji}e_{pq}=\sum_{j}\lambda_{p}S(e_{pj})\otimes e_{jq}$
    that $S(e_{pj})\in V_{\overline{\alpha}}$ for all $j,p$. This implies that $V_{\overline{\alpha}}=A_{\overline{\alpha}}$ for all $\alpha$ and hence $V=A$. By the way, from Proposition 1.3 in [VD-si1] we can conclude that the legs of $\Delta(h)$ are all of $A$, now in the sense of our paper here. Also Proposition 1.4 in [VD-si1] here says that $\Delta(h)(1\otimes a)=0$ only if $a=0$, and similarly $(a\otimes 1)\Delta(h)=0$ only if $a=0$. This we knew already because we have a multiplier Hopf algebra and $h\neq 0$.
    iii) Finally we come to Definition 1.5 of [VD-si1]. Since we know here already that
    $\Delta(h)(a\otimes 1)=\Delta(h(1\otimes S(a))\qquad\quad\text{and}\qquad\quad(1\otimes a)\Delta(h)=(S(a)\otimes 1)\Delta(h)$
    for all $a$, and because $S$ is bijective, we finally get that $\Delta(h)$ satisfies all the requirements and so it is a regular separability idempotent.
The maps $S:B\to C$ and $S^{\prime}:C\to B$ we have in Proposition 1.7 of [VD-si1] are precisely the antipode $S$ of $A$. If we look at the Proposition 1.8 in [VD-si1], we get here
$m(S\otimes\iota)(H(1\otimes e_{pq}))$ $=\sum_{i,j}\lambda_{i}S^{2}(e_{ij})e_{ji}e_{pq}$
$=\sum_{j}\lambda_{p}S^{2}(e_{pj})e_{jq}$
$=\sum_{j}\lambda_{j}e_{pj}e_{jq}=\sum_{j}\lambda_{j}e_{pq}$
and consequently $m(S\otimes\iota)(\Delta(h)(1\otimes a))=a$ for all $a$.
Similarly
$m((\iota\otimes S)((e_{pq}\otimes\iota)H)$ $=\sum_{i,j}\lambda_{i}\,e_{pq}S(e_{ij})S(e_{ji})$
$=\sum_{i,j}\lambda_{i}\,e_{pq}S(e_{ji}e_{ij})$
$=\sum_{i,j}\lambda_{i}\,e_{pq}S(e_{jj})$
$=\sum_{i}\lambda_{i}\,e_{pq}$
and consequently also $m(\iota\otimes S)\Delta(h))=a$ for all $a$.
The invariant integrals
Recall from Notation 2.10 the definition of the element $q_{\alpha}$ in $A_{\alpha}$. We have $q_{\alpha}=\\ (\omega_{\overline{\alpha}}\otimes\iota)\Delta(h)$ where $\omega_{\overline{\alpha}}$ is the usual trace on the component $A_{\overline{\alpha}}$. It is the unique element in $A_{\alpha}$ satisfying $aq_{\alpha}=q_{\alpha}S^{2}(a)$ for all $a\in A_{\alpha}$ and normalized such that $\omega_{\alpha}(q_{\alpha})=1$. It is a strictly positive element in $A_{\alpha}$.
We then get the following formula for the left integral on $A$.
2.18.
  • Theorem Let $\varphi$ be the left integral on $A$, normalized such that $\varphi(h)=1$. Then for each $\alpha$ we have $\varphi(a)=\omega_{\alpha}(aq_{\alpha}^{-1})$ when $a\in A_{\alpha}$.
  • Proof: Let $\alpha$ be an index in $I$ and consider the associated data we have used before. We found that
    $\sum_{i,j}\lambda_{i}S(e_{ij})\otimes e_{ji}=\Delta(h)(1_{\overline{\alpha}}\otimes 1_{\alpha})=\Delta(h)(1_{\overline{\alpha}}\otimes 1)$
    where $(e_{ij})$ are matrix units in $A_{\alpha}$, satisfying $q_{\alpha}=\sum_{i}\lambda_{i}e_{ii}$. Then
    $((\iota\otimes\omega_{\alpha}(\,\cdot\,q_{\alpha}^{-1}))\Delta(h))$ $=\sum_{i,j}\lambda_{i}S(e_{ij})\omega_{\alpha}(e_{ji}\lambda_{i}^{-1})$
    $=\sum_{i}\lambda_{i}S(e_{ii})\lambda_{i}^{-1}$
    $=S(1_{\alpha})=1_{\overline{\alpha}}.$
    This proves that $\omega_{\alpha}(aq_{\alpha}^{-1})=\varphi(a)$ for $a\in A_{\alpha}$.
We now obtain formulas for the other data, the modular automorphism $\sigma$, the modular element $\delta$ and the right integral $\psi$. We also verify some formulas relating these data.
2.19.
  • Proposition The $\alpha$-component of the modular element $\delta$ is $\omega_{\alpha}(q_{\alpha}^{-1})\,q_{\alpha}^{2}$.
  • Proof: Take an index $\alpha$ and consider the element $q_{\alpha}\in A_{\alpha}$ as defined in Notation 2.10. Then $aq_{\alpha}=q_{\alpha}S^{2}(a)$ and $S(q_{\alpha})S(a)=S^{3}(a)S(q_{\alpha})$ for all $a\in A_{\alpha}$. As $S(A_{\alpha})=A_{\overline{\alpha}}$ we have $aS(q_{\alpha}^{-1})=S(q_{\alpha}^{-1})S^{2}(a)$ for all $a\in A_{\overline{\alpha}}$. Therefore
    $q_{\overline{\alpha}}=\frac{1}{\omega_{\overline{\alpha}}(S(q_{\alpha}^{-1}))}S(q_{\alpha}^{-1})=\frac{1}{\omega_{\alpha}(q_{\alpha}^{-1})}S(q_{\alpha}^{-1}).$
    Then for $a\in A_{\overline{\alpha}}$,
    $\varphi(a)$ $=\omega_{\overline{\alpha}}(aq_{\overline{\alpha}}^{-1})=\omega_{\overline{\alpha}}(q_{\overline{\alpha}}^{-1}a)$
    $=\omega_{\alpha}(q_{\alpha}^{-1})\omega_{\overline{\alpha}}(S(q_{\alpha})a)$
    and we get
    $(\varphi\otimes\iota)\Delta(h)(1\otimes 1_{\alpha})$ $=\sum_{(h)}\omega_{\alpha}(q_{\alpha}^{-1})\omega_{\overline{\alpha}}(S(q_{\alpha})h_{(1)})h_{(2)}$
    $=\sum_{(h)}\omega_{\alpha}(q_{\alpha}^{-1})\omega_{\overline{\alpha}}(h_{(1)})q_{\alpha}h_{(2)}$
    $=\omega_{\alpha}(q_{\alpha}^{-1})\,q_{\alpha}^{2}.$
Recall that $a\delta=\delta S^{4}(a)$ for all $a$, see Proposition 1.16. And indeed, because $aq_{\alpha}=q_{\alpha}S^{2}(a)$, we have also $aq_{\alpha}^{2}=q_{\alpha}^{2}S^{4}(a)$ for $a\in A_{\alpha}$ so that the $\alpha$-component of $\delta$ has to be a scalar multiple of $q_{\alpha}^{2}$.
It is now easy to obtain the right integral $\psi$ because we have $\psi(a)=\varphi(S(a))$ since we have normalized such that $\psi(h)=\varphi(h)=1$. With this choice we get $\psi(a)=\varphi(a\delta)$ and so, for $a\in A_{\alpha}$ we find
$\psi(a)=\varphi(a\delta)=\omega_{\alpha}(a\delta q_{\alpha}^{-1})=\omega_{\alpha}(q_{\alpha}^{-1})\omega_{\alpha}(aq_{\alpha}).$
We immediately get the modular automorphisms from these formulas.
We have, for $a,b\in A_{\alpha}$,
$\varphi(ab)=\omega_{\alpha}(abq_{\alpha}^{-1})=\omega_{\alpha}(bq_{\alpha}^{-1}a)=\varphi(bq_{\alpha}^{-1}aq_{\alpha}).$
We see that $aq_{\alpha}=q_{\alpha}\sigma(a)$ and this is in agreement with $\sigma(a)=S^{2}(a)$.
Similarly, again for $a,b\in A_{\alpha}$,
$\psi(ab)=\omega_{\alpha}(q_{\alpha}^{-1})\omega_{\alpha}(abq_{\alpha})=\omega_{\alpha}(q_{\alpha}^{-1})\omega_{\alpha}(bq_{\alpha}a)=\psi(bq_{\alpha}aq_{\alpha}^{-1}).$
We see that $\sigma^{\prime}(a)=q_{\alpha}aq_{\alpha}^{-1}$ and so $aq_{\alpha}^{-1}=\sigma^{\prime}(a)q_{\alpha}^{-1}$. This is in agreement with $\sigma^{\prime}(a)=S^{-2}(a).$
3 . Duality between discrete and compact quantum groups
In this section, we will consider the dual of a discrete quantum group. We will follow the approach to duality as developed for general algebraic quantum groups in [VD-alg]. The dual will be a compact quantum group and we know that all compact quantum groups arise in this way. By looking at a compact quantum group as the dual of a discrete quantum group, we can obtain results about compact quantum groups in a different and easier way. We refer to the appendix for more comments on this matter.
The dual of a discrete quantum group
We start with a discrete quantum group as in Definition 2.2. So we have a direct sum $A$ of full matrix algebras $(A_{\alpha})_{\alpha\in I}$ and a coproduct $\Delta$ on $A$ as in Definition 2.1.
There are various ways to construct the dual, but in this case, the easiest way is by taking the reduced dual space as follows.
3.1.
  • Notation Let $B_{\alpha}$ be the dual $A^{\prime}_{\alpha}$ of $A_{\alpha}$ and use the pairing notation. So we write $\langle a,b\rangle$ for the value of an element $b$ from $B_{\alpha}$ in the point $a$ of $A_{\alpha}$. We use $B$ to denote the direct sum of the spaces $B_{\alpha}$. We have a pairing of $A$ with $B$ given by the formula $(a,b)\mapsto\sum_{\alpha\in I}\langle a_{\alpha},b_{\alpha}\rangle$ where now $a=(a_{\alpha})_{\alpha}$ and $b=(b_{\alpha})_{\alpha}$. The sum is well-defined because we only have finitely many terms that are non-zero. We consider $B$ as a subspace of the dual space $A^{\prime}$ of $A$.
Remark that for $b$ in the direct product of the spaces $B_{\alpha}$, we still can define $\langle a,b\rangle$ by this formula when $a\in A$ because also then we only have finitely many terms in the sum that are non-zero. Any element of $A^{\prime}$ is represented by an element in this direct product.
The space $B$ is made into an associative algebra by dualizing the coproduct of $A$. It is a -algebra. We make the following choices. We will motivate them later, see Remark 3.4.
In the following statement, $z^{-}$ denotes the complex conjugate of the complex number $z$.
3.2.
  • Proposition For $b,b^{\prime}\in B$ we use $\langle a,bb^{\prime}\rangle=\langle\Delta(a),b^{\prime}\otimes b\rangle$ for $a\in A$ to define the product $bb^{\prime}\in A^{\prime}$. For $b\in B$ we also define $b^{*}$ by $\langle a,b^{*}\rangle=\langle S(a^{*}),b\rangle^{-}$ for $a\in A$ These definitions make $B$ into a unital -algebra.
  • Proof: Because $b$ and $b^{\prime}$ are reduced functionals, so that they are supported on only finitely many components of $A$, this product is well-defined in the dual space $A^{\prime}$. It will again be in $B$. Associativity of this product is a consequence of coassociativity of the coproduct $\Delta$ on $A$. The counit $\varepsilon$ belongs to $B$ because it is a linear functional with support in the one-dimensional component $A_{e}$, spanned by the cointegral. It will be the unit in the algebra $B$.
    The map $b\mapsto b^{*}$ is clearly involutive as $S(S(a^{*})^{*})=a$ for all $a\in A$. And because $\Delta$ is a -homomorphism and $\Delta\circ S=\zeta(S\otimes S)\Delta$, we get $(bb^{\prime})^{*}={b^{\prime}}^{*}b^{*}$ for all $b,b^{\prime}\in B$. As before, we used $\zeta$ for the flip map. Therefore $B$ is a -algebra.
We also dualize the original product on $A$. The result is the following.
$3.$
  • Proposition The -algebra $B$ is a Hopf -algebra for the coproduct $\Delta$ defined by $\langle a\otimes a^{\prime},\Delta(b)\rangle=\langle aa^{\prime},b\rangle$ for $a,a^{\prime}\in A$ and $b\in B$.
  • Proof: i) When $b$ has its support in one component, say $A_{\alpha}$, then $\Delta(b)$ will have support in $A_{\alpha}\otimes A_{\alpha}$. Therefore $\Delta(b)\in B\otimes B$. Coassociativity of $\Delta$ on $B$ follows from associativity of the product on $A$. We have $\Delta(1)=1\otimes 1$ because $\varepsilon$ is a homomorphism on $A$. We get a homomorphism on $B$ because the coproduct $\Delta$ on $A$ is a homomorphism. Finally, because $S((aa^{\prime})^{*})=S(a^{\prime}{}^{*}a^{*})=S(a^{*})S(a^{\prime}{}^{*})$, we get that this coproduct is also a -homomorphism.
    ii) The counit $\varepsilon$ on $B$ is obtained by evaluation in $1$. The antipode $S$ on $B$ is given by the formula $\langle a,S(b)\rangle=\langle S^{-1}(a),b\rangle$ for $a\in A$.
In the formula for the antipode on the dual, the inverse appears as a result of our conventions. We give more details in the following remark.
3.4.
  • Remark i) In the theory of Hopf algebras, the conventions regarding duality are simple. The product on the dual is the dual of the coproduct and the coproduct on the dual is the dual of the product. This convention is completely symmetric. Unfortunately, in the operator algebra approach to quantum groups, things are a bit more complicated. Usually, the product on the dual is obtained by dualizing the coproduct (as for Hopf algebras) but the coproduct on the dual is obtained by flipping the coproduct coming from simply dualizing the product. This convention is of course not symmetric.
    ii) As mentioned already, discrete quantum groups first appeared as duals of compact quantum groups (within the operator algebra context) and in the more recent literature, this point of view is still quite common. This implies that the coproduct on the discrete quantum group is flipped. We are now looking reversely at a compact quantum group as dual to a discrete quantum group. Therefore, and in order to be in accordance with the existing literature on this subject, we have chosen to define the product on the dual $B$ of $A$ by using the flipped coproduct and thus we arrive at the formula $\langle a,bb^{\prime}\rangle=\langle\Delta(a),b^{\prime}\otimes b\rangle$ as in Proposition 3.2. On the other hand, the coproduct is defined as dual to the product giving the formula $\langle a\otimes a^{\prime},\Delta(b)\rangle=\langle aa^{\prime},b\rangle$ as in Proposition 3.3.
    iii) In this case, we also use $\Delta$ for the coproduct, $\varepsilon$ for the counit, $S$ for the antipode, etc. on $B$.
For convenience of the reader, below we collect the important formulas within this point of view.
3.5.
  • Remark For $a,a^{\prime}\in A$ and $b,b^{\prime}\in B$ we have
    $\langle a,bb^{\prime}\rangle$ $=\langle\Delta(a),b^{\prime}\otimes b\rangle$ (3.1)
    $\langle aa^{\prime},b\rangle$ $=\langle a\otimes a^{\prime},\Delta(b)\rangle$ (3.2)
    $\langle S(a),b\rangle$ $=\langle a,S^{-1}(b)\rangle$ (3.3)
    $\langle a,b^{*}\rangle$ $=\langle S(a^{*}),b\rangle^{-}$ (3.4)
    $\langle a^{*},b\rangle$ $=\langle a,S(b)^{*}\rangle^{-}.$ (3.5)
    Equation (3.5) follows from the Equations (3.3) and (3.4). Indeed we obtain
    $\langle a^{*},b\rangle$ $=\langle a^{*},b^{**}\rangle=\langle S(a),b^{*}\rangle^{-}$
    $=\langle a,S^{-1}(b^{*})\rangle^{-}=\langle a,S(b)^{*}\rangle^{-}$
    for all $a,b$.
Observe also the presence of the inverse of the antipode in Equation (3.3). This is one of the consequences of this convention. For the same reason, there is the lack of symmetry between (3.4) and (3.5). We do not get the usual formulas for a pairing of Hopf -algebras or multiplier Hopf -algebras (as e.g. in [Dr-VD]), but rather a modified form.
The integral on the dual
We now define the integral on the dual $(B,\Delta)$. Recall that $h$ is the cointegral in $A$.
3.6.
  • Proposition The map $b\mapsto\langle h,b\rangle$ is a positive linear functional on $B$. It is both left and right invariant.
  • Proof: Left and right invariance come from the fact that $ah=\varepsilon(a)h$ and $ha=\varepsilon(a)h$ for all $a$.
    To prove positivity take any $b\in B$. There exists an element $a\in A$ such that $b=\varphi_{A}(a\,\cdot\,)$ where $\varphi_{A}$ is the left integral on $A$ satisfying $\varphi_{A}(h)=1$. This is possible because $\varphi_{A}$ is faithful and $b$ is supported in only finitely many components of $A$. Then using the Sweedler notation,
    $\langle h,bb^{*}\rangle$ $=\langle\Delta(h),b^{*}\otimes b\rangle$
    $=\sum_{(h)}\langle h_{(1)},b^{*}\rangle\langle h_{(2)},b\rangle$
    $=\sum_{(h)}\langle h_{(1)},b^{*}\rangle\varphi_{A}(ah_{(2)})$
    $=\sum_{(h)}\langle S(a)h_{(1)},b^{*}\rangle\varphi_{A}(h_{(2)})$
    $=\langle S(a),b^{*}\rangle=\langle S(S(a)^{*}),b\rangle^{-}$
    $=\langle a^{*},b\rangle^{-}=\varphi_{A}(aa^{*})^{-}=\varphi_{A}(aa^{*})$
    and positivity follows from the positivity of $\varphi_{A}$.
So we see that the dual of the discrete quantum group (seen as a multiplier Hopf -algebra with positive integrals) is a Hopf -algebra with positive integrals. We also see from the above equation that the integral is faithful. This is in fact automatic (see e.g. [VD-alg]). We will encounter a similar result with a similar argument in Proposition 3.13 later in this section, as well in Proposition 4.6 the next section with a dual form.
3.7.
  • Proposition The modular automorphism $\sigma$ of $\varphi_{B}$ is given by
    $\langle a,\sigma(b)\rangle=\langle S^{-2}(a)\delta,b\rangle$
    where $a\in A$ and $b\in B$.
  • Proof: i) Let $d=\varphi(a\,\cdot\,)$ with $a\in A$. Then for all $c\in B$,
    $\varphi_{B}(dc)=\langle h,dc\rangle$ $=\sum_{(h)}\langle h_{(1)},c\rangle\langle h_{(2)},d\rangle$
    $=\sum_{(h)}\langle h_{(1)},c\rangle\varphi(ah_{(2)})$
    $=\sum_{(h)}\langle S(a)h_{(1)},c\rangle\varphi(h_{(2)})$
    $=\langle S(a),c\rangle.$
    ii) On the other hand, for all $b\in B$,
    $\varphi_{B}(bd)=\langle h,bd\rangle$ $=\sum_{(h)}\langle h_{(1)},d\rangle\langle h_{(2)},b\rangle$
    $=\sum_{(h)}\varphi(ah_{(1)})\langle h_{(2)},b\rangle)$
    $=\sum_{(h)}\varphi(h_{(1)})\langle S^{-1}(a)h_{(2)},b\rangle)$
    $=\langle S^{-1}(a)\delta,b\rangle.$
    iii) It follows that $\langle S(a),\sigma(b)\rangle=\langle S^{-1}(a)\delta,b\rangle$ for all $a$ and hence
    $\langle a,\sigma(b)\rangle=\langle S^{-2}(a)\delta,b\rangle$
    for all $a\in A$ and $b\in B$.
Let us make a remark about this formula for the modular automorphism $\sigma$ of the integral $\varphi_{B}$ on $B$.
3.8.
  • Remark i) The formulas for our convention are obtained by taking the opposite product on the algebra $B$ while the coproduct remains the same. Consequently, the left and right integrals remain the same, as well as the modular element. On the other hand, the modular automorphisms should be replaced by their inverses.
    ii) For the usual duality between algebraic quantum groups $A$ and $B$ we have the formula for $\delta_{B}$ given by $\langle a,\delta_{B}\rangle=\varepsilon(\sigma^{-1}(a))$. We use the extension of the pairing to $A\times M(B)$. Because in our case, the modular automorphism $\sigma$ is the same as the square of the antipode (see Proposition 1.13, we will have $\langle a,\delta_{B}\rangle=\varepsilon(a)$ and this means that $\delta_{B}=1$. This is in agreement with the fact that the left integral is also right invariant.
    iii) Again for the usual duality we have the formula $\langle a,\sigma(b)\rangle=\langle S^{2}(a)\delta^{-1},b\rangle$. Then for the inverse of the modular automorphism, we get $\langle S^{-2}(a)\delta,\sigma^{-1}(b)\rangle$. This is in agreement with the formula we got in the previous proposition because, passing to our conventions, the antipode on $A$ remains the same, as well as the modular element, while the modular automorphism on $B$ has to be replaced by its inverse.
Formulas of this type are found at various places in the literature. One possible reference is [De-VD1] where they are obtained in the more general case of algebraic quantum hypergroups. They are also found in [VD-part1].
If we let $B$ act on the GNS-representation space associated with this integral, we can complete it to a C-algebra. Also the coproduct can be extended. One of the key features that makes this possible is the fact that the -algebra is represented by bounded operators in the GNS-space associated with the positive integral. This is not true for general positive linear functionals on a -algebra. This result is already found in [Ku-VD], see also [DC-VD] and the more recent approach in [VD-part2].
In what follows we will sometimes denote the integral on the dual compact quantum group $B$, that we obtained in Proposition 3.6, by $\varphi_{B}$ and we will use $\varphi_{A}$ and $\psi_{A}$ for the left and the right integrals on the original discrete quantum group $A$. If no confusion is possible, we will drop these subscripts.
Corepresentations of the dual
As one can expect, there is a one-to-one correspondence between corepresentations of the dual $(B,\Delta)$ and representations of the discrete quantum group $(A,\Delta)$. We make this precise and examine it in detail.
Let $n$ be given and consider an element $u\in B\otimes M_{n}(\mathbb{C})$. Associate a linear map $\pi$ from $A$ to $M_{n}(\mathbb{C})$ by $\pi(a)=(\omega_{a}\otimes\iota)u$ where $\omega_{a}=\langle a,\,\cdot\,\rangle$. In what follows, we use the leg-numbering notation as explained in the introduction.
3.9.
  • Proposition The map $\pi$ is a homomorphism if and only if $(\Delta\otimes\iota)u=u_{13}u_{23}$. It is a non-degenerate homomorphism if and only if $u$ is invertible. In that case the inverse of $u$ is $(S\otimes\iota)u$. It is a -representation if and only if $u$ is a unitary element in the -algebra $B\otimes M_{n}(\mathbb{C})$.
  • Proof: i) Given $a,a^{\prime}\in A$ we have
    $\pi(aa^{\prime})$ $=(\omega_{aa^{\prime}}\otimes\iota)u=((\omega_{a}\otimes\omega_{a^{\prime}}\otimes\iota)((\Delta\otimes\iota)u)$
    $\pi(a)\pi(a^{\prime})$ $=((\omega_{a}\otimes\iota)u)((\omega_{a^{\prime}}\otimes\iota)u)=(\omega_{a}\otimes\omega_{a^{\prime}}\otimes\iota)(u_{13}u_{23}).$
    If $\pi(aa^{\prime})=\pi(a)\pi(a^{\prime})$ for all $a,a^{\prime}$, it follows that $(\Delta\otimes\iota)u=u_{13}u_{23}$. Clearly also the converse is true.
    ii) Assume now that $\pi$ is a homomorphism such that $(\Delta\otimes\iota)u=u_{13}u_{23}$ holds. If we apply the counit $\varepsilon$ of $B$ on the first factor of the equation $(\Delta\otimes\iota)u=u_{13}u_{23}$, we see that $u=eu$ where $e=(\varepsilon\otimes\iota)u$. Similarly if we apply it on the second factor. Then we get $ue=u$. If we apply $m(S\otimes\iota)$ on the equation $(\Delta\otimes\iota)u=u_{13}u_{23}$ we find
    $((S\otimes\iota)u)u=(\varepsilon\otimes\iota)u=e$
    and similarly, if we apply $m(\iota\otimes S)$ we get $(u(S\otimes\iota)u)=e$. We see that $u$ is invertible if and only if $e=1$. This happens if and only if the representation $\pi$ is non-degenerate.
    iii) Finally, if $u$ is a unitary we find
    $\pi(a)^{*}$ $=((\omega_{a}\otimes\iota)u)^{*}=(\overline{\omega_{a}}\otimes\iota)(u^{*})$
    $=(\overline{\omega_{a}}\otimes\iota)((S\otimes 1)u).$
    But we have
    $\overline{\omega_{a}}(S(b))=\omega_{a}(S(b)^{*})^{-}=\langle a,S(b)^{*}\rangle^{-}=\langle a^{*},b\rangle.$
    This means that $\overline{\omega_{a}}\circ S=\omega_{a^{*}}$ and so $\pi(a)^{*}=\pi(a^{*})$. It is also easy to see that conversely, when $\pi(a)^{*}=\pi(a^{*})$ for all $a$, we will have that $u^{*}=(S\otimes\iota)u$ so that $u$ is a unitary element in $B\otimes M_{n}(\mathbb{C})$.
We also have a converse of the previous result.
3.10.
  • Proposition For every $n$-dimensional representation $\pi$ of $A$, there is an element $u\in B\otimes M_{n}(\mathbb{C})$ such that $\pi(a)=(\omega_{a}\otimes\iota)u$ for all $a\in A$.
  • Proof: Let $\pi$ be an $n$-dimensional representation of $A$. Define linear functionals $u_{ij}$ on $A$ by $\langle a,u_{ij}\rangle=\pi(a)_{ij}$. We claim that these functionals must be elements in $B$. Then we put $u=\sum_{ij}u_{ij}\otimes e_{ij}$ where $e_{ij}$ are matrix units. We clearly will have
    $(\omega(a\,\cdot\,)\otimes\iota)u=\sum_{ij}\langle a,u_{ij}\rangle e_{ij}=\sum_{ij}\pi(a)_{ij}e_{ij}=\pi(a).$
    To prove the claim, observe that for only finitely many indices $\alpha\in I$ we can have $\pi(e_{\alpha})\neq 0$. So the functionals $u_{ij}$ will be supported on finitely many components and hence they belong to $B$.
This in the end shows that indeed, there is a one-to-one correspondence between $n$-dimensional representations of $A$ and elements in elements $u\in B\otimes M_{n}(\mathbb{C})$ satisfying $(\Delta\otimes\iota)u=u_{13}u_{23}$.
Let us now have a look at the more general case of a possibly infinite-dimensional corepresentation with coefficients in the C-algebra completion $\overline{B}$ of $B$ and not just in $B$. We now only look at unitary corepresentations. So in this case $u$ is in the multiplier algebra $M(\overline{B}\otimes\mathcal{K}(\mathcal{H}))$ where $\mathcal{K}(\mathcal{H})$ is the C-algebra of compact operators on a Hilbert space $\mathcal{H}$. The C-tensor product is taken. It is still required that $(\Delta\otimes\iota)u=u_{13}u_{23}$, this time in the multiplier algebra $M(\overline{B}\otimes\overline{B}\otimes\mathcal{K}(\mathcal{H}))$.
We can again associate a -representation of $A$ on the Hilbert space $\mathcal{H}$ using the same formula, but we will need the following lemma to do this.
3.11.
  • Lemma For all $a\in A$ the linear map $b\mapsto\langle a,b\rangle$ is continuous on $B$ and therefore can be extended to the C-algebra $\overline{B}$.
  • Proof: We have
    $a=(\iota\otimes\varphi)(\Delta(h)(a\otimes 1))=(\iota\otimes\varphi)(\Delta(h)(1\otimes S(a))$
    so that $\langle a,b\rangle=\varphi_{B}(b^{\prime}b)$ with $b^{\prime}=\varphi(\,\cdot\,S(a))$. Because $b\mapsto\varphi_{B}(b^{\prime}b)$ is continuous, this proves the result.
Then we can show the following.
3.12.
  • Proposition For a unitary corepresentation $u\in M(\overline{B}\otimes\mathcal{K}(\mathcal{H}))$ we can define a non-degenerate -representation $\pi$ of $A$ on $\mathcal{H}$ by $\pi(a)=(\omega_{a}\otimes\iota)u$ where as before $\omega_{a}=\langle a,\,\cdot\,\rangle$, now extended to $\overline{B}$.
  • Proof: Remark that $\pi(a)$ is defined as a multiplier of $\mathcal{K}(\mathcal{H})$ and so $\pi(a)$ is defined in $\mathcal{B}(\mathcal{H})$. The rest of the proof is essentially the same as for finite-dimensional corepresentations. $\square$
If we take any central projection $e$ of $A$ and if we cut down $\pi$ by $e$, we see that the resulting representation lives on a finite number of components of $A$. It follows that $u(1\otimes\pi(e))$ is actually in the algebraic tensor product $B\otimes\mathcal{B}(\pi(e)\mathcal{H})$. Therefore, $u$ is a multiplier of the algebraic tensor product $B\otimes\mathcal{K}_{0}(\mathcal{H})$ where $\mathcal{K}_{0}(\mathcal{H})$ denotes the -subalgebra of operators living on these finite-dimensional subspaces. In particular, if $\mathcal{H}$ is finite-dimensional, then $u\in B\otimes\mathcal{B}(\mathcal{H})$ and we get a corepresentation as we defined it earlier.
The left regular corepresentation
Next we look at the left regular corepresentation of $B$ and show that it gives the GNS-representation of $A$ associated with the left integral $\varphi_{A}$. The unitary realizing the equivalence is the Fourier transform.
Consider the GNS-representation of $B$ associated with the integral and denote by $\Lambda_{B}:B\to\mathcal{H}_{B}$ the associated canonical map. We let $B$ act directly on the Hibert space $\mathcal{H}_{B}$ so that $\Lambda_{B}(bb^{\prime})=b\Lambda_{B}(b^{\prime})$ for all $b,b^{\prime}\in B$. Similarly, take the GNS-representation of $A$ for the left integral $\varphi$ on $A$ and denote the canonical map by $\Lambda_{A}:A\to\mathcal{H}_{A}$.
3.13.
  • Proposition With the above notations, there is a unitary operator $F:\mathcal{H}_{A}\to\mathcal{H}_{B}$ given by $F\Lambda_{A}(a)=\Lambda_{B}(b)$ were $a\in A$ and $b=\varphi_{A}(S^{-1}(\,\cdot\,)a)$. If $\pi$ is the -representation of $A$ on $\mathcal{H}_{B}$ determined by the left regular corepresentation of $B$, then $F^{*}\pi(\,\cdot\,)F$ is the GNS representation of $A$ on $\mathcal{H}_{A}$.
  • Proof: i) First, we will show that $\varphi_{B}(b^{*}b)=\varphi_{A}(a^{*}a)$ if $a\in A$ and $b=\varphi_{A}(S^{-1}(\,\cdot\,)a)$. This is a modified form of the formula found in the proof of Proposition 3.6. The argument is essentially the same. We have
    $\varphi_{B}(b^{*}b)$ $=\langle h,b^{*}b\rangle=\langle\Delta(h),b\otimes b^{*}\rangle$
    $=\sum_{(h)}\langle h_{(1)},b\rangle\langle h_{(2)},b^{*}\rangle$
    $=\sum_{(h)}\varphi_{A}(S^{-1}(h_{(1)})a)\langle h_{(2)},b^{*}\rangle$
    $=\sum_{(h)}\psi_{A}(S(a)h_{(1)})\langle h_{(2)},b^{*}\rangle$
    $=\sum_{(h)}\psi_{A}(h_{(1)})\langle ah_{(2)},b^{*}\rangle$
    $=\langle a,b^{*}\rangle=\langle S(a^{*}),b\rangle^{-}$
    $=\varphi_{A}(a^{*}a)^{-}=\varphi_{A}(a^{*}a).$
    This implies the existence of the unitary operator $F:\mathcal{H}_{A}\to\mathcal{H}_{B}$ given by $F\Lambda_{A}(a)=\Lambda_{B}(b)$.
    ii) The left regular corepresentation $u$ is defined by
    $u^{*}(\eta\otimes\Lambda_{B}(b))=\sum_{(b)}b_{(1)}\eta\otimes\Lambda_{B}(b_{(2)}),$
    see the papers on locally compact quantum groups, mentioned in the introduction. Because $u^{*}=(S\otimes\iota)u$ we get
    $u(\eta\otimes\Lambda_{B}(b))=\sum_{(b)}S^{-1}(b_{(1)})\eta\otimes\Lambda_{B}(b_{(2)}).$
    Therefore, with $\pi$ as in Proposition 3.12 we find
    $\pi(a^{\prime})\Lambda_{B}(b)=\sum_{(b)}\langle a^{\prime},S^{-1}(b_{(1)})\rangle\Lambda_{B}(b_{(2)})$
    when $a^{\prime}$ is another element in $A$. Now
    $\sum_{(b)}\langle a^{\prime},S^{-1}(b_{(1)})\rangle$ $\langle x,b_{(2)}\rangle=\langle S(a^{\prime})x,b\rangle$
    $=\varphi_{A}(S^{-1}(S(a^{\prime})x)a)=\varphi_{A}(S^{-1}(x)a^{\prime}a)$
    and therefore
    $\sum_{(b)}\langle a^{\prime},S^{-1}(b_{(1)})\rangle\Lambda_{B}(b_{(2)})=F\Lambda_{A}(a^{\prime}a)=Fa^{\prime}\Lambda_{A}(a).$
    Because $\Lambda_{B}(b)=F\Lambda_{A}(a)$, we get indeed $F^{*}\pi(a^{\prime})F=a^{\prime}$ on $\mathcal{H}_{A}$.
More formulas
Finally, we look at some important formula we find in the theory of compact quantum groups and see what the corresponding results are for the discrete quantum groups.
The orthogonality relations of irreducible unitary corepresentations yield an operator $Q$ in $M(A)$ by the formula
$\varphi_{B}(u^{\alpha}_{ip}(u^{\beta}_{jq})^{*})=\delta_{\alpha\beta}\delta_{ij}Q^{\alpha}_{qp}.$
In the following formula, we find that this element is nothing else but (a scalar multiple of) $\delta^{-\frac{1}{2}}$ where now $\delta$ is the modular element of the discrete quantum group, relating the left and the right integrals.
3.14.
  • Proposition Choose an index $\alpha$ and denote by $v$ the associated unitary corepresentation of $B$. The
    $\varphi_{B}(v_{ij}(v_{ik})^{*})=(q_{\alpha})_{jk}.$
  • Proof: We know that $\varphi_{A}(aa^{*})=\varphi_{B}(bb^{*})$ if $b=\varphi_{A}(a\,\cdot\,)$ (see Proposition 3.6). So we have to find the elements $a_{ij}\in A_{\alpha}$ such that $v_{ij}=\varphi_{A}(a_{ij}\,\cdot\,)$. This means that we must have
    $\varphi_{A}(a_{ij}x)=\langle x,v_{ij}\rangle=x_{ij}$
    when $x\in A_{\alpha}$. So
    $x_{ij}=\omega_{\alpha}(a_{ij}xq_{\alpha}^{-1})=\omega_{\alpha}(q_{\alpha}^{-1}a_{ij}x).$
    Therefore $q_{\alpha}^{-1}a_{ij}=e_{ji}$ where we have matrix units in $A_{\alpha}$. Then
    $\varphi_{B}(v_{ij}v_{ik}^{*})$ $=\varphi_{A}(a_{ij}a_{ik}^{*})$
    $=\omega_{\alpha}(q_{\alpha}e_{ji}e_{ik}q_{\alpha}q_{\alpha}^{-1})$
    $=\omega_{\alpha}(q_{\alpha}e_{jk})=(q_{\alpha})_{j}\delta_{jk}.$
4 . Discrete and compact groups
In this section, we will discuss two special cases to illustrate the results of the previous sections. We begin with the motivating case of a discrete quantum group arising from a discrete group. Most of this is trivial but we include it for completeness. There are also some consequences of the choice we made as in Remark 3.5 in the previous section.
We also look at the dual of a compact group. In [VD-suq2] we treat a more advanced example in greater detail.
The discrete quantum group arising from a discrete group
Recall Example 2.3 from Section 2. In that example, $G$ is any group. A discrete quantum space (as in the beginning of that section) is associated. The index set is identified with the set $G$ and let $n(p)=1$ for all $p$ in $G$. So every component $A_{p}$ is equal to $\mathbb{C}$. The direct sum $A$ of these components is nothing else but the algebra $K(G)$ of all complex functions on $G$ with finite support and pointwise operations. A coproduct on $A$ is defined by $\Delta(f)(p,q)=f(p,q)$. Then the pair $(A,\Delta)$ is a discrete quantum group as in Definition 2.2.
Here the algebra $A$ is abelian. We also have the converse result.
4.1.
  • Proposition If $(A,\Delta)$ is a discrete quantum group and if $A$ is abelian, then it is the discrete quantum group associated to a group as above.
  • Proof: Now identify the set $G$ with the index set $I$ of the discrete quantum group $(A,\Delta)$. For all pairs $p,q\in G$ we consider the homomorphism $f\mapsto\Delta(f)(p,q)$ form $A$ to $\mathbb{C}$. It is given by evaluation in a point of $G$ that we denote by $pq$. This gives an associative product on $G$ because $\Delta$ is coassociative. There is an identity $e$ for the multiplication defined by $f(e)=\varepsilon(f)$. And every element $p\in G$ has an inverse $p^{-1}$ given by $f(p^{-1})=S(f)(p)$ for all $f\in K(G)$.
In this case the cointegral $h$ is the function $\delta_{e}$ with value $1$ in $e$ and 0 everywhere else. For $\Delta(h)$ we can write
$\Delta(h)=\sum_{p}\delta_{p}\otimes\delta_{p^{-1}},$
where just like for $\delta_{e}$, the function $\delta_{p}$ takes the value $1$ in the point $p$ and 0 everywhere else.
The left integral $\varphi_{A}$ is given by $\varphi(f)=\sum_{p}f(p)$. It is also a right integral. It is straightforward to verify left and right invariance of $\varphi_{A}$. Indeed we have
$(\iota\otimes\varphi_{A})\Delta(h)$ $=\sum_{p}\delta_{p}\varphi_{A}(\delta_{p^{-1}})=\sum_{p}\delta_{p}=1,$
$(\varphi_{A}\otimes\iota)\Delta(h)$ $=\sum_{p}\varphi_{A}(\delta_{p})\delta_{p^{-1}}=\sum_{p}\delta_{p^{-1}}=1.$
These equations hold in the multiplier algebra of $K(G)$.
As the algebra is abelian, these functionals are traces and the modular automorphisms are trivial. Also the modular element $\delta$ is trivial because the left and right integrals coincide. The square $S^{2}$ of the antipode is the identity map.
Let us have a quick look at the formula in Proposition 2.11. For the element $q$, defined in Notation 2.10, we get here
$q=(\omega\otimes\iota)\Delta(h)=\sum_{p}\omega(\delta_{p})\delta_{p^{-1}}=\sum_{p}\delta_{p^{-1}}=1.$
Indeed, the normalized trace on the components are given $\delta_{p}\mapsto 1$ for all $p$. So as expected, we get $q=1$ in this case.
We now pass to the results of Section 3 for this example.
As before, let $G$ be a group and $(K(G),\Delta)$ the associated discrete quantum group. We also consider the Hopf -algebra $(\mathbb{C}G,\Delta)$. Here $\mathbb{C}G$ is the group algebra and $\Delta$ the coproduct on $\mathbb{C}G$ given by $\Delta(\lambda_{p})=\lambda_{p}\otimes\lambda_{p}$. We use $p\mapsto\lambda_{p}$ for the canonical embedding of the group in the group algebra.
4.2.
  • Proposition Define a linear map $\gamma$ from $\mathbb{C}G$ to the dual of $K(G)$ by
    $\gamma(\lambda_{p})=\varphi_{A}(\,\cdot\,\delta_{p^{-1}}).$
    Then $\gamma$ is an isomorphism from $(\mathbb{C}G,\Delta)$ to the dual of $(K(G),\Delta)$.
  • Proof: It is clear that the map $\gamma$ is a bijective map from the space $\mathbb{C}G$ to the underlying space of the dual of the discrete quantum group $(K(G),\Delta)$.
    Consider the pairing of $K(G)$ with its dual. Then we have
    $\langle f,\gamma(\lambda_{p})\rangle=\varphi(f\delta_{p^{-1}})=f(p^{-1})$
    for all $f\in K(G)$ and $p\in G$.
    i) Given $f\in K(G)$ and $p,q\in G$ we have
    $\langle f,\gamma(\lambda_{pq})\rangle$ $=f((pq)^{-1})=f(q^{-1}p^{-1})$
    $=\Delta(f)(q^{-1},p^{-1})$
    $=\langle\Delta(f),\gamma(\lambda_{q})\otimes\gamma(\lambda_{p})\rangle$
    $=\langle f,\gamma(\lambda_{p})\gamma(\lambda_{q})\rangle\$
    Remember the convention about the product on the dual algebra (cf. Equation (3.1) in the previous section). So we get $\gamma(\lambda_{p}\lambda_{q})=\gamma(\lambda_{p})\gamma(\lambda_{q})$ and $\gamma$ is a homomorphism from the group algebra $\mathbb{C}G$ to the dual of $K(G)$.
    ii) Let $f\in K(G)$ and $p\in G$. Then
    $\langle f,\gamma(\lambda_{p}^{*})\rangle$ $=\langle f,\gamma(\lambda_{p^{-1}})\rangle=f(p),$
    $\langle f,\gamma(\lambda_{p})^{*}\rangle$ $=\langle S(f^{*}),\gamma(\lambda_{p})\rangle^{-}$
    $=(S(\overline{f})(p^{-1})^{-}=f(p)$
    and we see that $\gamma(\lambda_{p}^{*})=\gamma(\lambda_{p})^{*}$. Here we have used Equation (3.4) from the previous section. Therefore $\gamma$ is a -isomorphism from $\mathbb{C}G$ to the dual of $K(G)$.
    iii) Given $f,g\in K(G)$ and $p\in G$ we have (using Equation (3.2),
    $\langle fg,\gamma(\lambda_{p})\rangle$ $=(fg)(p^{-1})=f(p^{-1})g(p^{-1})$
    $=\langle f\otimes g,\gamma(\lambda_{p})\otimes\gamma(\lambda_{p})\rangle$
    $=\langle f\otimes g,(\gamma\otimes\gamma)\Delta(\lambda_{p})\rangle.$
    This means that $\Delta(\gamma(\lambda_{p}))=(\gamma\otimes\gamma)\Delta(\lambda_{p})$. This completes the proof.
The above result gives the pairing of $K(G)$ and $\mathbb{C}G$ as in the following remark.
4.3.
  • Remark Consider the -isomorphism $\gamma$ from $\mathbb{C}G$ to $\widehat{K(G)}$ given in the previous proposition. It induces a pairing of $K(G)$ with $\mathbb{C}G$ by the formula
    $\langle f,\lambda_{p}\rangle=\langle f,\gamma(\lambda_{p})\rangle=f(p^{-1}).$
    Again because of our conventions, this is not the usual pairing of these two algebras.
Let us also consider the dual integral. The right integral $\psi_{B}$ on the dual is defined by $\psi_{B}(b)=\varepsilon(f)$ when $b=\varphi_{A}(\,\cdot\,f)$. Hence we get
$\psi_{B}(\gamma(\lambda_{p}))=\varepsilon(\delta_{p^{-1}})=\delta(e,p^{-1})=\delta(e,p)$
and indeed $\psi_{B}\circ\gamma$ is the integral on the group algebra $\mathbb{C}G$.
The dual of the group algebra $\mathbb{C}G$
In this item, we start with a group $G$ and its group algebra $\mathbb{C}G$. We also use $B$ for this algebra. We consider the coproduct $\Delta$ defined by $\Delta(\lambda_{p})=\lambda_{p}\otimes\lambda_{p}$. Recall that $p\mapsto\lambda_{p}$ is the canonical embedding of $G$ in $\mathbb{C}G$. This gives a Hopf -algebra with integrals. The left integral $\varphi_{B}$ is also a right integral and $\varphi_{B}(\lambda_{p})=0$ except for $p=e$. Recall that we use $e$ for the identity of the group $G$. If we normalize the left integral $\varphi_{B}$ by letting $\varphi_{B}(\lambda_{e})=1$, we get a positive integral.
Indeed, we have for $b=\sum_{p}b(p)\lambda_{p}$ that
$\varphi_{B}(b^{*}b)=\sum_{p,q}\overline{b(p)}b(q)\varphi(\lambda_{p^{-1}q})=\sum_{p}\overline{b(p)}b(p).$
[…]
In what follows, we will use $A$ for the algebra $\widehat{B}$ and consider the pairing notation $(a,b)\mapsto\langle a,b\rangle$ where $a\in A$ and $b\in B$. With these conventions, the previous statement gives, more precisely, the following result.
$4.$
  • Proposition Define $\rho:K(G)\to A$ by $\rho(\delta_{p})=\varphi_{B}(\,\cdot\,\lambda_{p})$ for all $p\in G$. Then $\rho$ is an isomorphism of the discrete quantum group $(K(G),\Delta)$ to the dual of the Hopf -algebra $(\mathbb{C}G,\Delta)$.
  • Proof:
    It is clear that $\rho$ is an isomorphism of vector spaces. For any $f\in K(G)$ and $p\in G$ we have
    $\langle\rho(f),\lambda_{p}\rangle=\sum_{q}f(q)\varphi_{B}(\lambda_{p}\lambda_{q})=\sum_{q}f(q)\varphi_{B}(\lambda_{pq})=f(p^{-1}).$
    ii) For any pair of functions $f,g\in K(G)$ we have, for all $p\in G$ that
    $\langle\rho(fg),\lambda_{p}\rangle$ $=(fg)(p^{-1})=f(p^{-1})g(p^{-1})$
    $=\langle\rho(f),\lambda_{p}\rangle\langle\rho(g),\lambda_{p}\rangle$
    $=\langle\rho(f)\otimes\rho(g),\lambda_{p}\otimes\lambda_{p}\rangle$
    $=\langle\rho(f)\otimes\rho(g),\Delta(\lambda_{p})\rangle$
    $=\langle\rho(f)\rho(g),\lambda_{p}\rangle$
    and so $\rho(fg)=\rho(f)\rho(g)$ so that $\rho$ is a homomorphism. We have used Equation (3.2).
    ii) We also have, using Equation (3.5),
    $\langle\rho(f^{*}),\lambda_{p}\rangle$ $=f^{*}(p^{-1})=\overline{f}(p^{-1})$
    $\langle\rho(f)^{*},\lambda_{p}\rangle$ $=\langle f,S(\lambda_{p})^{*}\rangle^{-}$
    $=\langle f,\lambda_{p}\rangle^{-}=\overline{f}(p^{-1})$
    and we see that because $\rho(f^{*})=\rho(f)^{*}$ . So we have that $\gamma$ is a -homomorphism
    iii) Finally we show that $\Delta(\rho(f))=(\rho\otimes\rho)\Delta(f)$ for all $f\in K(G)$. To show this, take $p,q\in G$. On the one hand we have, using Equation (3.1),
    $\langle\Delta(\rho(f)),\lambda_{p}\otimes\lambda_{q}\rangle=\langle\rho(f),\lambda_{q}\lambda_{p}\rangle=\langle\rho(f),\lambda_{qp}\rangle=f(p^{-1}q^{-1})$
    while on the other hand
    $\langle(\rho\otimes\rho)\Delta(f),\lambda_{p}\otimes\lambda_{q}\rangle=\Delta(f)(p^{-1},q^{-1})=f(p^{-1}q^{-1}).$
    This completes the proof.
This -isomorphism induces the same pairing between $K(G)$ and $\mathbb{C}G$ we have in the previous item. Indeed
$\langle f,\lambda_{p}\rangle=\langle\rho(f),\lambda_{p}\rangle=f(p^{-1}).$
In this equation, the left hand side is the pairing of $K(G)$ and $\mathbb{C}G$ as it induced by this isomorphism from the pairing of the dual of $\mathbb{C}G$ with $\mathbb{C}G$ in the middle expression. The last equality is by the definition of $\rho$ as we have seen in the beginning of the previous proof.
The dual of a compact quantum group
Next we start with a more general setting than in the previous item. We now consider any Hopf -algebra with positive integrals. These are the so-called -algebraic quantum groups of compact type, or shortly, as we will call them here, the compact quantum groups.
So in what follows $(B,\Delta)$ is a Hopf -algebra with a positive integral $\varphi_{B}$. From the positivity of the integral, we must have that $\varphi_{B}(1)\geq 0$ and we can assume that $\varphi_{B}(1)=1$. It also follows that $\varphi_{B}$ is not only left invariant, but also right invariant.
We will now show that the dual $(A,\Delta)$ of $(B,\Delta)$, in the sense of duality of algebraic quantum groups, is a discrete quantum group as defined in Definition 2.2 of Section 1. We will however work within the conventions used in Section 3, cf. the formulas in Remark 3.5.
First we have the existence of a cointegral.
4.5.
  • Proposition The integral $\varphi_{B}$ on $B$ is a left cointegral in $A$. It is a self-adjoint idempotent.
  • Proof: We define $h$ in $A$ by $\varphi_{B}$. This belongs to the dual of $B$ because $B$ has an identity.
    i) Let $a\in A$. Then for all $y\in B$, using that $\varphi_{B}$ is right invariant, we have
    $\langle ah,y\rangle$ $=\langle a\otimes h,\Delta(y),\rangle$
    $=\langle a,(\iota\otimes\varphi_{B})\Delta(y)\rangle$
    $=\varphi_{B}(y)\langle a,1\rangle=\langle h,y\rangle\,\varepsilon(a)$
    and we see that $ah=\varepsilon(a)h$. Similarly, because $\varphi_{B}$ is also right invariant, we get $ah=\varepsilon(a)h$.
    ii) For the adjoint $h^{*}$ of $h$ in $B$ we get
    $\langle h^{*},y\rangle=\langle h,S(x)^{*}\rangle^{-}=\varphi_{B}(S(y)^{*}))^{-}=\varphi_{B}(S(y))=\varphi_{B}(y)$
    and $h^{*}=h$. We have used that $\varphi_{B}\circ S=\varphi_{B}$ here and that $\varphi_{B}$ is self-adjoint.
    iii) For $h^{2}$ we get $\varepsilon(h)h$ and $\varepsilon(h)=\langle 1,h\rangle=\varphi_{B}(1)=1.$
To conclude that the dual $(A,\Delta)$ of $(A,\Delta)$ is a discrete quantum group, we use Theorem 1.10 that we already proved in Section 1. For this, we need that the dual $A$ is an operator algebra. This will follow from the positivity of the integral. For this we can use the general result that the dual has a positive integral if the original $(B,\Delta)$ has a positive integral.
We include the result below and we keep using our conventions from Section 3. For these conventions, as the coproduct on the dual $A$ is flipped (cf. Equation (3.1), now the left integral $\varphi_{A}$ is given by the formula $\varphi_{A}(a)=\varepsilon(b)$ when $a=\varphi_{B}(\,\cdot\,b)$ with $b\in A$.
4.6.
  • Proposition For any $a\in A$ of the form $\varphi_{B}(\,\cdot\,b)$ we get $\varphi_{A}(a^{*}a)=\varphi_{B}(b^{*}b)$.
  • Proof: For all $y$ in $B$ we have
    $\langle a^{*}a,y\rangle$ $=\langle a^{*}\otimes a,\Delta(y)\rangle$
    $=\sum_{(y)}\langle a^{*},y_{(1)}\rangle\langle a,y_{(2)}\rangle$
    $=\sum_{(y)}\langle a^{*},y_{(1)}\rangle\varphi_{B}(y_{(2)}b)$
    $=\sum_{(b)}\langle a^{*},S^{-1}(b_{(1)})\rangle\varphi_{B}(yb_{(2)})$
    and it follows that
    $a^{*}a=\sum_{(b)}\langle a^{*},S^{-1}(b_{(1)})\rangle\varphi_{B}(\,\cdot\,b_{(2)}).$
    Then
    $\varphi_{A}(a^{*}a)$ $=\sum_{(b)}\langle a^{*},S^{-1}(b_{(1)})\rangle\varepsilon(b_{(2)})$
    $=\langle a^{*},S^{-1}(b)\rangle=\langle a,b^{*}\rangle^{-}$
    $=\varphi_{B}(b^{*}b)^{-}=\varphi_{B}(b^{*}b).$
We have use the Equations (3.2) and (3.5) from Section 3, as well as the formula
$S((\iota\otimes\varphi_{B})(\Delta(y)(1\otimes b)))=(\iota\otimes\varphi_{B})((1\otimes y)\Delta(b)).$
This formula is true for any left integral on an algebraic quantum group. It is the dual form of the equality
$(1\otimes a)\Delta(h)=(S(a)\otimes 1)\Delta(h)$
as we explain further in Remark 4.8.
Compare the above argument with the proof of Proposition 3.6. There we get the positivity of the integral on $B$ from the positivity of the integral on $A$. Both results are essentially based on the same formula. This formula is like the Plancherel formula for locally compact abelian groups, see more details about this in [VD-part2].
Now we can complete the proof of the following result.
4.7.
  • Theorem The dual of a Hopf -algebra with a positive integral is a discrete quantum group.
  • Proof: Take a Hopf -algebra $B$ with a positive integral $\varphi_{B}$ as before. Let $A$ be the dual multiplier Hopf -algebra in the sense of duality of algebraic quantum groups. From Proposition 4.5 we know that $A$ has a cointegral that is a self-adjoint idempotent. From Proposition 4.6 we know that $A$ again has a positive integral. The $A$ is an operator algebra and from from Theorem 1.10 we get that $A$ is a direct sum of finite-dimensional matrix algebras. Then it is a discrete quantum group by Definition 2.2
We are using that a multiplier Hopf -algebra with a positive integral is an operator algebra. In this case, because of the structure of $A$, this is an easy result. In general, it uses that the algebra is represented by bounded operator in the GNS representation of $A$ induced by the positive functional. See e.g. [VD-part2] for details.
4.8.
  • Remark Take $x,y$ in $B$ and $a\in A$. We then have on the one hand
    $\langle(1\otimes a)\Delta(h),x\otimes y\rangle$ $=\sum_{(h)}\langle h_{(1)},x\rangle\langle ah_{(2)},y\rangle$
    $=\sum_{(h),(y)}\langle h_{(1)},x\rangle\langle a,y_{(1)}\rangle\langle h_{(2)},y_{(2)}\rangle$
    $=\sum_{(y)}\langle h,y_{(2)}x\rangle\langle a,y_{(1)}\rangle$
    $=\langle a\otimes h,\Delta(y)(1\otimes x)\rangle$
    This is the pairing of $a$ with $(\iota\otimes\varphi_{B})(\Delta(y)(1\otimes x))$. On the other hand
    $\langle(S(a)\otimes 1)\Delta(h),x\otimes y\rangle$ $=\sum_{(h)}\langle S(a)h_{(1)},x\rangle\langle h_{(2)},y\rangle$
    $=\sum_{(h),(x)}\langle S(a),x_{(1)}\rangle\langle h_{(1)},x_{(2)}\rangle\langle h_{(2)},y\rangle$
    $=\sum_{(x)}\langle S(a),x_{(1)}\rangle\langle h,yx_{(2)}\rangle$
    $=\langle S(a)\otimes h,(1\otimes y)\Delta(x)\rangle.$
    This is the pairing of $S(a)$ with $(\iota\otimes\varphi_{B})((1\otimes y)\Delta(x))$. We see that the equality
    $(1\otimes a)\Delta(h)=(S(a)\otimes 1)\Delta(h)$
    in $A\otimes A$ is nothing else but the adjoint of the formula
    $(\iota\otimes\varphi_{B})(\Delta(y)(1\otimes x))=S^{-1}((\iota\otimes\varphi_{B})((1\otimes y)\Delta(x)))$
    in $B$. We have used the formulas we have in Remark 3.5.
5 . Conclusion and further research
In this paper, we have given a treatment of discrete quantum groups. It is different from the original one in [VD-discrete]. The study of discrete quantum groups started before other more general developments in the theory of locally compact groups. We felt an update was needed with references to these newer developments.
An important role is played by the element $\Delta(h)$. It is a separability idempotent in the sense of [VD-si1] and many of its properties are in fact coming from this fact.
We consider the dual of a discrete quantum group in the sense of duality of algebraic quantum groups as in [VD-alg]. It is a compact quantum group and properties of the compact quantum group follow within this point of view of properties of discrete quantum groups.
In another paper on the subject [VD-suq2], the discrete quantum group of the quantized enveloping algebra of the Lie algebra of $SU(2)$ is studied in this setting. It illustrates many of the general results as described in this paper. It could be a challenging project to study more of such concrete examples in this framework.
A review of the more recent results on discrete quantum groups, obtained by various authors but with the approach of considering discrete quantum groups as duals of compact quantum groups also seems a desirable and interesting project.
Appendix. Different approaches to discrete quantum groups
In this appendix, we will discuss and compare different approaches to discrete quantum groups and we will place our treatment in this context. We will also mention differences in conventions used within these different approaches.
Discrete quantum groups as duals of compact quantum groups
Compact quantum groups in the operator algebra framework are introduced and studied by Woronowicz. First he only considered a restricted class of such compact quantum groups in [Wo2] while later, he studied them from a more general point of view [Wo3]. This last paper only appeared in 1995, but it was available as a preprint many years before that. In this paper, we have used that a compact quantum group is in fact just a Hopf -algebra with a positive integral.
It seems appropriate to start this discussion with the work of Podleś and Woronowicz as they were the first to consider discrete quantum groups in an operator algebraic setting. We look at the Sections 2 and 3 in their paper on the quantum deformation of the Lorentz group [Po-Wo].
In Section 2, they first review the theory of compact matrix quantum groups as developed in [Wo2]. The underlying C-algebra is denoted by $A_{c}$ and we will denote the coproduct by $\Delta_{c}$. Then they use the results about corepresentations of $(A_{c},\Delta_{c})$ to construct the underlying algebra $A_{d}$ of the dual discrete quantum group. It is defined as a C-direct sum of matrix algebras. Corepresentation theory of $A_{c}$ is also used to construct the ’left regular corepresentation’ $u$ in the multiplier algebra $M(A_{d}\otimes A_{c})$ of the C-tensor product of $A_{d}$ with $A_{c}$.
The first main result in [Po-Wo] is Theorem 3.1. The coproduct on the dual $A_{d}$ is constructed, together with a counit and an antipode satisfying the standard equalities. The coproduct is defined as a non-degenerate -homomorphism $\Delta_{d}:A_{d}\to M(A_{d}\otimes A_{d})$ by the formula $(\Delta_{d}\otimes\iota)u=u_{23}u_{13}$. This means that they work with the ’flipped’ coproduct on the dual, i.e. the usual convention in the operator algebra approach. The counit is a -homomorphism from $A_{d}$ to $\mathbb{C}$ (and so everywhere defined and bounded). The antipode however is not defined everywhere but only on the algebraic direct sum, denoted by $A_{d0}$, of the components of $A_{d}$ and maps this dense -subalgebra into itself.
In their Theorem 3.3, they obtain the left and the right integrals, defined on the subalgebra $A_{d0}$. They also find that the modular element, relating the left with the right integral, is equal to (a scalar multiple of) $F^{2}$ where the element $F$ in the multiplier algebra of the subalgebra that was already constructed within the theory of compact matrix quantum groups, using the orthogonality relations of the matrix elements of the irreducible corepresentations.
Finally in Theorem 3.4, they essentially show that the dual of $A_{d}$ gives again $A_{c}$.
So what is done by Podleś and Woronowicz in these two sections of [Po-Wo] can be summarized as follows. They define a discrete quantum group as a pair of a C-algebra AA with a coproduct $\Delta$ on $A$, where $A$ is a direct sum of matrix algebras and so that furthermore there is a counit and an antipode. They associate to each compact quantum matrix group a dual discrete quantum group and they obtain further results about this discrete quantum group (such as the existence of the integrals) using this fact and various results about compact quantum groups and their corepresentation theory.
Comparison with the multiplier Hopf algebra approach
In [VD-discrete], a discrete quantum group is defined as a multiplier Hopf -algebra $(A,\Delta)$ where $A$ is a (algebraic) direct sum of matrix algebras (see Definition 2.3 in [VD-discrete]). A multiplier Hopf -algebra is defined in [VD-mha], Definitions 2.3 and 2.4.
This definition is not very different from the one given by Podleś and Woronowiz in the sense that it takes little effort to pass from one to the other. Indeed, if we start e.g. with a multiplier Hopf -algebra, the general theory allows to construct a counit and an antipode (see Section 3 and 4 in [VD-mha]). Then, as the algebra is a direct sum of matrix algebras, it is not hard to extend the coproduct and the counit to the C-direct sum. This takes us to a discrete quantum group as introduced (implicitly) in Theorem 3.1 of [Po-Wo].
Conversely, if we start with a discrete quantum group as in [Po-Wo], we simply restrict the coproduct to the algebraic direct sum of the components and we obtain a discrete quantum group in the sense of [VD-discrete]. The counit and the antipode are used to show that the basic axioms for a multiplier Hopf -algebra (see [VD-mha] and/or [VD-discrete]) are fulfilled. It is worthwhile noticing that the axioms for multiplier Hopf algebras can also be formulated in terms of the existence of a counit and an antipode (as for Hopf algebras) and then the two definitions come even closer together.
So the difference between the two approaches does not really lie in the definition of what is a discrete quantum group, but rather in other aspects. We will discuss these now.
The first, probably the main difference is that in [VD-discrete] the existence of the invariant integrals is proven from the axioms. In [Po-Wo], the existence of the integrals on the discrete quantum group is obtained given the fact that it is the dual of a compact quantum group. However, if you want to show that every discrete quantum group is the dual of a compact one, you need the integrals to do that.
Along the same lines, by treating the discrete quantum groups ad duals of compact quantum groups, also properties of discrete quantum groups are proven using results about compact quantum groups. In the other, independent approach, all results about discrete quantum groups are obtained from the axioms.
The second difference, of another nature but maybe equally important, is the fact that we consider the duality between compact and discrete quantum groups as a special case of the duality for algebraic quantum groups. In this theory, the general case turns out to be simpler than the special case. If e.g. the general construction of the dual (as for algebraic quantum groups) is applied to the special cases of a compact or a discrete quantum group, it is easy to see that the result is a discrete, respectively a compact quantum group. The arguments are simpler than the ones given in [Po-Wo].
This leads in general, also at other places where recently other aspects of compact and discrete quantum groups are studied, to simplifications and a better understanding. This should not come as a surprise because it is already the case in the approach to discrete quantum groups itself as treated in [Po-Wo].
Compact quantum groups as duals of discrete quantum groups
As one can study discrete quantum groups as duals of compact quantum groups, similarly it is possible to study compact quantum groups as duals of discrete quantum groups. This is possible because on the one hand, the discrete quantum groups, as well as the compact quantum groups can be studied on their own, while on the other hand, it is shown in each of the two cases that the dual of one type is an object of the other type. However, it must be said in order to get the full picture, one needs to develop each of the theories, up to a certain extend, separately. One has e.g. to prove the existence of the integrals in each of the two cases.
Starting with a discrete quantum group and from there, studying compact quantum groups is indeed possible (and is somewhat easier than the other, more common way). We show this in these notes. But this only makes sense if you know already that any compact quantum group is the dual of a discrete quantum group. And as mentioned already, this result requires to develop the compact quantum groups independently. But the same is true for the other direction.
Taken abstraction of this (and of the historical context), people who want to work with compact and discrete quantum groups can consider any of the two approaches. To a great extend, it is a matter of taste which you take as a starting point (although we are convinced that starting with discrete quantum groups is a bit easier). However, for a full understanding, we feel it is worthwile to study the two theories also as independent entities.